### v1

# Raising the Bar of AI-generated Image Detection with CLIP

Abstract
Aim of this work is to explore the potential of pre-trained vision-language models (VLMs) for universal detection of AI-generated images. We develop a lightweight detection strategy based on CLIP features and study its performance in a wide variety of challenging scenarios. We find that, unlike previous belief, it is neither necessary nor convenient to use a large domain-specific dataset for training. On the contrary, by using only a handful of example images from a single generative model, a CLIP-based detector exhibits a surprising generalization ability and high robustness across several different architectures, including recent commercial tools such as Dalle-3, Midjourney v5, and Firefly. We match the SoTA on in-distribution data, and improve largely above it in terms of generalization to out-of-distribution data (+6% in terms of AUC) and robustness to impaired/laundered data (+13%). Our project is available at https://grip-unina.github.io/ClipBased-SyntheticImageDetection/
1 Introduction
Synthetic images have by now left the research laboratories and are flooding the real world. The latest versions of popular image editing tools, such as Adobe Photoshop and Microsoft Paint, come with easy-to-use AI-powered generative tools that allow even novice users to edit and generate visual content at will. Thanks to diffusion-based generative models, it is not only the quality of the images generated that has improved enormously, but also the flexibility in using these tools. By issuing a few textual commands, it is possible to easily obtain the desired image. While this represents a great opportunity for visual art applications, it is also the paradise of disinformation professionals who can design their attacks with unprecedented power and flexibility [8, 5, 25]. And, of course, a nightmare for all those trying to combat the spread of fake news and orchestrated disinformation campaigns: journalists, fact-checkers, law enforcement agencies, governments. Therefore, there is high demand for automatic tools that help establishing the authenticity of a media asset.
Figure 1: Area Under ROC Curve (AUC %) on unseen synthetic generators ($x$-axis) and on post-processed data ($y$-axis). The first number measures the generalization ability of the detector, the second its robustness to possible impairments. Circle area proportional to training set size. The performance is measured over 18 different synthetic models. Our CLIP-based detector largely outperforms all SOTA methods with very limited training data.
It is well known that each synthesis model leaves its own peculiar traces in all generated images, subtle traces that give rise to so called artificial fingerprints, that can be exploited for forensic analyses [49, 77]. This was first observed for methods based on generative adversarial networks (GAN) [79, 22] but holds with obvious differences for all generation approaches, including the more recent diffusion models (DM) [13]. One of the main challenges for current forensic detectors is the ability to generalize to new and unseen generative methods. Indeed, the in-distribution (ID) scenario, with perfectly aligned training and test set, is rarely met in practice. Test images are often generated by new approaches, unseen architectures, or even known architectures re-trained under different conditions, all different flavors of the out-of-distribution (OOD) scenario. In addition, most of the times, images are downloaded from some social networks, where they undergo transformations such as compression and resizing (even multiple times), processes that tend to wash out the tiny traces so precious for detection.
In [71], it is shown that training diversity and intense augmentation are crucial for generalization. A ResNet-50 detector is trained on a single but very diverse dataset of about 360k ProGAN images [40]. Results on OOD images generated by other GAN-based generators turned out to be surprisingly good. The authors themselves, however, suggest that structural similarities between GAN-based generators are the likely reason for such good performance. Indeed, results are not equally good on images generated by recent diffusion-based methods, which present somewhat different generation artifacts [14].
Generalizing to OOD data is a major issue in deep learning and has been the object of intense research in latest years. In this context, the advent of large pre-trained vision-language models has brought about a number of new solutions and exciting results. These models have been shown to be excellent zero-shot and few-shot learners in many diverse applications, such as image classification [59, 78], detection [26, 53] and segmentation [74, 81]. Recently, there have been attempts to exploit the power of VLMs to detect synthetic images [66, 2, 56]. For example, in [56] CLIP (Contrastive Language-Image Pre-Training) [59] is used as a pre-trained feature extractor, and a classifier is then trained on the 360k-image ProGAN dataset proposed in [71], with the same augmentation strategy. With respect to [71] based on Resnet50, the performance improves significantly, especially on images generated by diffusion models, showing a very good generalization ability.
In this work we explore in depth the potential of CLIP for image forensics, propose new decision strategies, and carry out an extensive experimental analysis in challenging real-world scenarios involving a large number of generative models. We find that CLIP-based methods show impressive generalization ability and very good robustness, resulting in large performance improvements compared to the state of the art. To obtain such results we avoid any intensive training on domain-specific data, which may introduce unwanted biases and undermine the descriptive power of CLIP features. Instead, we rely on a small number of paired real/fake images with the same textual description and use their CLIP features to model the decision space. Synthetic images come from a single generator but work well with all other tested data sources. The top performance is obtained with just 1000 to 10000 paired images. Moreover, only minor decays are observed when this number reduces to 100 or even 10 (see Fig. 1).
In summary, the key contributions of this work are as follows:
  • We show that CLIP features achieve excellent generalization: by exploiting only a handful of examples, not even belonging to the generator under test, the performances are comparable to those of intensively trained solutions.
  • We carry out a large set of experiments on diverse synthetic generators and in very challenging conditions and are able to achieve the best performance (on average). Our experiments make clear that the features extracted are at least partially orthogonal to those explored by previous methods mostly oriented to extract low-level features.
2 Related work
There is an extensive literature on synthetic image detection, mostly concerning images created by generative adversarial networks (GANs) and, more recently, by diffusion models (DMs). Some methods look for visible errors, such as asymmetries in faces, wrong perspective, odd shadows [51, 28, 29]. However, with the rapid progress in image synthesis technology, this kind of problems are solved quickly and appear more and more rarely in modern generation methods. Hence, we will focus on methods that exploit “invisible” forensic traces, working either in the spatial or the frequency domain. Then we will briefly review the most recent approaches that exploit multimodal features.
Spatial domain methods. Despite their high visual quality, synthetic images carry with them distinctive traces of the generation process that enable detection and even model identification [49, 77]. In fact, each generation model inserts a sort of digital fingerprint into all the images created, which depends on its architectural and training details. This fingerprint can be easily estimated, Given an adequate number of images generated by the target model, typically a few hundred, it is sufficient to average their noise residuals, extracted through simple denoisers [49] or more sophisticated deep learning-based methods [67, 47]. Detectors based on digital fingerprints can be regarded as “few-shot”, as they can deal with new models based on just a few example images. Other few-shot solutions have been proposed lately in this field [15, 50, 21, 36]. However, they all require some prior knowledge of the target models, even if limited to a few images, so they fit a strictly in-distribution (ID) scenario. In contrast, our work aims to generalize to new and previously unseen models, and therefore fits into an out-of-distribution (OOD) scenario.
Good generalization is a key requirement of synthetic image detectors. Towards this end, it is important to increase diversity in training, which can be pursued by suitable augmentation, by including a large number of different categories [71], or even by model ensembling [48]. Other studies suggest to work on local patches [9] or combine global spatial information with local features [38]. In [33] it is shown that to preserve the subtle high-frequency forensic traces, one should avoid any downsampling in the first layers of the network. With the same aim, [69] proposes to work on noise residuals rather than the original images, by extracting the gradients with a pre-trained CNN. In [12] a systematic study is carried out on transferable forensic features that allows easier generalization.
The above investigations and proposals, however, only consider GAN-based generators. Some very recent works extend the analyses to latest generation approaches. [14] shows that detectors designed for GAN images have a hard time generalizing to DM images, especially in the presence of common post-processing steps such as compression and resizing. Furthermore, they are unable to select an adequate decision threshold without the help of some calibration data from the model under test. On the other hand, [24] shows that it is possible to achieve decent performance by continuously re-training the detector on images from new generators, as long as the latter share some architectural components with the known ones. In [72], inspired to previous work on GAN images [1], a detector specifically tailored to DM images is proposed. Images are projected in a latent space and reconstructed with known models to study their features. While promising, the approach refers to specific architectures, hence an ID scenario, and has been tested on a limited set of diffusion generative models. In the present work, instead, we do not make any assumptions on the test data and include a large variety of both GAN and DM generators including OOD images, and realistic high-resolution images generated by recent diffusion models.
Frequency domain methods. GAN-image artifacts are more easily spotted in the frequency domain and are clearly visible in the artificial fingerprint spectra [22, 31, 23]. In fact, they are caused by the up-sampling operations used in the generator, which give rise to regular spatial patterns and strong peaks in the Fourier domain. Interestingly, the awareness of such weakness has prompted the design of new architectures, like StyleGAN3 [43], which explicitly avoid aliasing and reduce the above mentioned peaks. In addition to such obvious artifacts, the spectral content of GAN images and real images is also known to differ significantly in the medium and high frequencies. Likewise, clear differences in the radial and angular spectral distributions of the DM and real images were observed [13, 75]. Furthermore, it is worth observing that DM images also may present spectral peaks (see Fig. 5).
Of course, frequency domain traces can be used to train simple detectors for ID images. Even more interestingly, some authors [79, 37] replicate and modify the synthesis process of known generators by introducing a series of small architectural changes in order to learn to manage also OOD test images. A major problem with spectral traces is that they can be easily weakened by laundering operations, both innocent (e.g., image resizing) and malicious counterforensic methods [20, 11]. Likewise, attackers can easily embed artificial spectral traces in a real image to make it appear as synthetic. This is a more general problem of low-level traces, which can be attacked in both ways and through a number of ingenious methods. We will show that our approach is indeed immune to such attacks and can handle a variety of distribution shift.
Exploiting multimodal features. The introduction of large language models has elicited intense work in the vision community on trying to exploit large models trained using both images and text [80, 78]. In image forensics, however, traditionally oriented to the analysis of low-level features, there has been very little work in this direction. Only a few papers [66, 2, 56] have tried exploiting pre-trained visual models for the detection of AI-generated images. All of them rely on variants of CLIP, however only in [56] it is pursued generalization to OOD data, by nearest neighbor and linear probing in the CLIP feature space. To this end, a large dataset of fake and real images is used to train the classifier. In this work we make a further step in this direction and show that better performance can be achieved even using much less data.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 2: Examples of synthetic images from generators used in our experiments. From left to right, Top: GLIDE [55], Latent Diffusion [60], DALL·E 2 [60]. Middle: Stable Diffusion 1.3, Stable Diffusion 1.4, Stable Diffusion 2.1 [62]. Bottom: Stable Diffusion XL [58], Adobe Firefly [30], DALL·E 3 [6].
3 Datasets and metrics
The main objective of this work is to study the generalization ability of CLIP-based and conventional methods, that is, assess how the detectors behave on data never seen before, of different provenance, and possibly impaired by various forms of post-processing. Therefore, we consider a very large test dataset and try to avoid any possible biases in the experimental validation.
Tab. 1 lists all the generation models used in testing. They are grouped in three families: GAN-based (ProGAN, StyleGAN2, StyleGAN3, StyleGAN-T, GigaGAN), DM-based (Score-SDE, ADM, GLIDE, eDiff-I, Latent and Stable Diffusion, DiT, DeepFloyd-IF, Stable Diffusion XL) and commercial tools (DALL·E 2, DALL·E 3, Midjourney v5 and Adobe Firefly), which includes some recent text-driven methods available on-line, whose architecture is not always disclosed. These models cover the different generation modalities commonly adopted in this field: unconditional (u), conditional (c), and text-to-image (t), reported in second column of the table.
The last column reports the resolution of the images in the dataset. The central columns, instead, indicate which datasets of real images have been considered to carry out the detection experiments. These have been carefully selected to avoid all foreseeable biases. For example, when testing on synthetic images generated by ProGAN, we use as real counterparts images drawn from the LSUN dataset on which the ProGAN model was trained. Therefore, if LSUN images shared all the same defect (say, a red ”LSUN” label in a corner) the ProGAN images should spot the same defect, which could not be used to spoil detection performance. This hilarious example should not hide the importance of the problem, as biases are way too common in experiments and not always so easy to avoid. In other cases, for example when text-driven commercial tools with undisclosed training details are used, we relied on a dataset recently proposed in [4]. Pristine and synthetic images with similar semantic content are built by extracting textual descriptions from a real dataset (RAISE [16]) and then using them as prompt for generating synthetic ones. Overall, we selected $32,000$ real and fake images for testing, which will form the dataset used in all forthcoming experiments. To consider a more realistic scenario, we also consider datasets with post-processing operations that simulate images shared over social networks. In particular, we apply random cropping, resizing and JPEG compression at different quality levels. More details on the dataset can be found in the supplemental material, here we only show a few examples in Fig. 2.
To measure performance, we consider three metrics. The area under the receiver-operating curve (AUC), and the average precision (AP), equal to the area under the precision-recall curve, are both threshold-independent. The Accuracy instead, given by the number of correct predictions divided by the number of tested images, depends on the decision threshold. We use always a fixed threshold equal to 0.5, which fits a realistic scenario with no prior information on the data under test to carry out proper calibration.
Generator
modality
LSUN [76]
FFHQ [41]
ImageNet [18]
COCO [46]
LAION [64]
RAISE [16]
Resolution
ProGAN [40] u $256^{2}$
StyleGAN2 [42] u $256^{2}$-$1024^{2}$
StyleGAN3 [43] u $256^{2}$-$1024^{2}$
StyleGAN-T [63] t $512^{2}$
GigaGAN [39] c,t $256^{2}$,$512^{2}$
Score-SDE [68] u $256^{2}$
ADM [19] u,c $256^{2}$
GLIDE [55] t $256^{2}$
eDiff-I [3] t $256^{2}$,$1024^{2}$
Latent Diff. [61] u,c,t $256^{2}$
Stable Diff. [62] t $256^{2}$-$768^{2}$
DiT [57] c $256^{2}$,$512^{2}$
DeepFloyd-IF [44] t $1024^{2}$
Stable Diff. XL [58] t $1024^{2}$
DALL·E 2 [60] t $1024^{2}$
DALL·E 3 [6] t $1024^{2}$
Midjourney V5 [52] t $1024^{2}$-$1100^{2}$
Adobe Firefly [30] t $2032^{2}$-$2048^{2}$
Table 1: Image generators used in our experiments: GAN-based, DM-based, and commercial tools with models not always disclosed. For unconditional (u) and conditional (c) models, to prevent biases, the real images used in test come from the same datasets used to train the generator. Image resolution varies from 256$\times$256 to 2048$\times$2048.
4 CLIP for synthetic image detection
We propose a simple procedure to tell apart real images from synthetic images based on features extracted from the image encoder of CLIP ViT L/14. The design of the detector consists of the following four steps:
  1. 1.
    collect $N$ real images $\{R_{1},\ldots,R_{N}\}$ with the corresponding captions $\{t_{1},\ldots,t_{N}\}$;
  2. 2.
    use the captions to feed a text-driven image generator, $G(\cdot)$, so as to obtain $N$ synthetic images $\{F_{1},\ldots,F_{N}\}$ with $F_{i}=G(t_{i})$. We have now $N$ real / fake pairs that share the same textual description;
  3. 3.
    feed CLIP with the $N$ real and $N$ fake images and collect the corresponding feature vectors $\{{\bf r}_{1},\ldots,{\bf r}_{N}\}$ and $\{{\bf f}_{1},\ldots,{\bf f}_{N}\}$ extracted at the output of the next-to-last layer, with ${\bf r}_{i}={\rm CLIP}(R_{i})$ and ${\bf f}_{i}={\rm CLIP}(F_{i})$;
  4. 4.
    use the $N$+$N$ feature vectors to design a linear SVM classifier.
If the real images have low-quality associated captions, or no caption at all, these can be easily generated by a dedicated tool such as BLIP [45]. Real and fake images are coupled to avoid possible semantic biases, which proves beneficial when the number of images is very small, e.g., $N=10$. We chose to extract features from the next-to-last layer rather than the last layer based on the results of preliminary experiments and, in the same way, we selected SVM among a set of candidate simple classifiers. All these preliminary experiments and their results are described in the supplemental material.
In the following, we analyze the performance of the proposed detector as a function of the main design choices and experimental conditions.
4.1 Influence of the reference set size
We begin with investigating the role of the most important parameter of the proposed detector, the number of real / synthetic samples $N$ in the reference set, that in this experiment includes COCO real images and Latent Diffusion generated ones. Fig. 3 shows results for $N$ varying from 10 to 100k in log scale. The results are presented in terms of AUC, AP, and Accuracy. When $N$ is less than 10k, we perform several runs with different sets of images and report the average metrics, e.g. for $N$ equals to 1k, we show the average on 10 runs.
Let us focus first on the top row, showing results on images that have been neither re-compressed nor resized. In this simpler scenario, AUC and AP are extremely high, always above 85% for all known families of synthetic images (GAN, Diffusion) and somewhat worse for images generated by commercial tools. Especially remarkable is the fact that very good results are obtained already with as few as 10+10 reference images, showing this lightweight solution to be fully viable. The performance improves with $N$, growing by up to 5-10 percent points at the $N=10k$ plateau. We observe, in passing, that AUC and AP curves are very similar, as always happens when they approach 100%, hence we drop the latter for brevity in further experiments.
Figure 3: Performance of the CLIP-based detector as a function of the number of real and synthetic images in the reference set. We show AUC, AP and Accuracy on the original dataset (Top) and on post-processed images that simulate a realistic scenario (Bottom).
In the rightmost figure, we see that unlike AUC and AP, accuracy does not always improve with increasing $N$. This is because AUC and AP are integral threshold-independent metrics. They tell us how well a method could do with a perfectly calibrated threshold. On the contrary, accuracy depends critically on the selected threshold which, lacking any prior knowledge, is set to 0.5. For Latent and GAN, 0.5 is a good threshold, and accuracy improves as the classifier does. Instead, for Diffusion and Commercial, the optimal threshold grows farther from 0.5 as $N$ increases and the accuracy decreases significantly. So there is a trade-off between optimality and robustness in our truly OOD scenario, where no calibration data exist (it is very unlikely that we know exactly which is the generator of the data under test). On the other hand, if we had data from the target class, even just 10 images, we could use them to design and ad hoc classifier.
The bottom row of the figure shows results for images that have been compressed and/or resized. As expected, performance degrades somewhat in this scenario, but continues to be very good, with AUC between 75% and 90% and accuracy between 65% and 80% at $N$=10k. These are excellent results, much beyond the current state of the art, as we will show later. In this case, using only 10+10 images seems not to be advisable but the performance is near-optimal already at $N$=100.
Datasets
augm.
GAN
Diffusion
Commerc.
AVG
family
family
tools
COCO $+$ Latent
93.9
93.3
81.7
89.6
COCO $+$ Latent
84.8
91.0
87.6
87.8
COCO $+$ ProGAN
95.0
90.0
65.6
83.5
COCO $+$ ProGAN
92.3
90.1
82.3
88.2
LSUN $+$ Latent
87.0
82.7
65.6
78.5
LSUN $+$ Latent
86.8
79.9
80.3
82.4
LSUN $+$ ProGAN
93.8
79.2
44.7
72.6
LSUN $+$ ProGAN
95.7
82.7
67.5
82.0
Table 2: Performance of the CLIP-based detector for various real/fake combinations, with and without augmentation.
4.2 Influence of the reference set content
We discovered that not only the size but also the quality of the reference set images impacts heavily on performance. This probably does not need proof if we think about a very small reference set, like 10+10 images, where low quality or low diversity would spell catastrophe. However, something similar happens also when a larger reference set is used. Tab. 2 shows the results in terms of AUC obtained using various combination of real and synthetic images in the reference set, that is, (Real, Synth) $\in\{{\rm COCO,LSUN}\}\times\{{\rm Latent,ProGAN}\}$, with and without augmentation. In all cases, the reference set has size 10000+10000. Results are clearly affected by the specific combination. Indeed, there is a marked impairment when the LSUN dataset is used instead of COCO to draw the real samples and also, to a lesser extent, when ProGAN is used instead of Latent for the synthetic images. We conjecture that both LSUN and ProGAN are not diverse enough to describe adequately the decision domain. Moreover, LSUN has several polarizations not included in COCO, the images of LSUN represent only few well-defined categories, have all the same size, and are compressed most of them with the same quality factor. In Tab. 2, we see that data augmentation only marginally reduces the effects of the used dataset.
Figure 4: Performance of the CLIP-based detector as a function of the pre-training. We show AUC, AP and Accuracy on post-processed images for models pre-trained on LAION-400M (0.4B images), LAION (2B) and CommonPool (12.8B).
4.3 Influence of pre-training
We experimented also with various versions of CLIP ViT L/14 [35], trained on different datasets, observing significant changes in performance. Fig. 4 shows, the AUC, AP and Accuracy averaged on all families of models as a function of the pre-training dataset. Each curve refers to a different value of $N$ and we consider only the case of post processed images for brevity. For both metrics and all values of $N$ there is a steady increase in performance as larger and larger datasets are used for pre-training, from LAION-400M (0.4B images) [64], to LAION (2B) [65], to CommonPool (12.8B) [32]. Overall, from smallest to largest dataset there is a ten point improvement! This comes to confirm, if ever necessary, the importance of pre-training a large VLM on the largest possible collection of varied images to acquire greater descriptive power for all kinds of applications.
4.4 Beyond low-level forensic traces
As already highlighted in the Introduction, forensic classifiers typically rely on subtle low-level artifacts. These can be seen more clearly in the frequency domain, where they appear as strong peaks. In this subsection we report on some initial analysis aimed at understanding whether the CLIP-based detector also relies on this type of traces or it exploits different information. To this end, we consider the high resolution 1024$\times$1024-pixel images generated by Stable Diffusion XL and resize them to 256$\times$256. Fig. 5 shows, on the right, the high and low-resolution images (top) together with their spectra (bottom). We observe that after resizing, the spectral artifacts have completely disappeared. In fact, these traces are caused by the aliasing effect of the upsampling filters, which can be removed by proper decimation. We run our detector on 1000 original and resized synthetic images, obtaining in both cases the same True Positive Rate (TPR) of $43\%$ (note that the True Negative Rate (TNR) for these images is $97\%$).
We also carry out a complementary experiment where we take real images and pass them through the generative architecture of Stable Diffusion which acts as an autoencoder. Therefore, the output (Fig. 5 right) is visually identical to the input but its Fourier spectrum now shows strong peaks. Even so, the CLIP-based detector is not affected by the presence of such traces and the TNR, $97\%$, is almost the same as with unprocessed images. We find this last experiment to be particularly relevant. In fact, it is very important to take into account malicious operations that can be carried out with the aim of reducing the credibility of current forensic detectors. Of course, several forms of attack can be applied, but this one seems to be especially simple and disruptive, such to easily fool detectors that rely exclusively on low-level traces. Based on this experiment, it seems that the proposed CLIP-based detector looks for different types of traces. Therefore, besides its stand-alone performance, it could be used to complement and enhance the performance of other detectors. A larger set of experiments can be found in the supplementary.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: Top (from left to right): a high resolution synthetic image generated by Stable Diffusion XL [58] and its 4$\times$ decimated version; a real image and the corresponding image obtained at the output of Stable Diffusion [62] used as an autoencoder. Bottom: Fourier spectra of the noise residuals for images shown on the top. A suitable decimation removes Fourier peaks in synthetic images, while passing a real image through an autoencoder creates new peaks.
GAN family Diffusion family AVG
Pro
Style
Style
Style
Giga
Score
ADM GLIDE
Latent
Stable
DeepFl.
Ediff-I DiT SDXL
Method
real/synth. training
size (k)
GAN
GAN2
GAN3
GANT
GAN
SDE
Diff.
Diff.
IF
Wang et al.
LSUN / ProGAN
360 / 360
100.
96.5
89.0
98.9
66.6
32.9
64.3
48.5
59.2
41.5
78.0
64.9
58.6
54.3
68.1
PatchFor.
CelebA,FF / various
84 / 272
92.3
84.5
87.9
91.2
64.7
83.3
74.8
96.2
78.1
62.4
62.7
78.7
83.1
68.4
79.2
Grag. et al.
LSUN / ProGAN
360 / 360
100.
99.8
98.3
98.8
82.8
92.1
74.7
62.8
91.9
52.5
69.9
69.6
65.3
58.0
79.8
Liu et al.
LSUN / ProGAN
360 / 360
100. 99.8
97.8
98.5
98.2 95.4
82.5
76.5
97.6
77.4
72.2
98.7
88.0
31.1
86.7
Corvi et al.
COCO,LSUN / Latent
180 / 180
79.4
73.7
38.4
97.1
39.9
65.0
80.7
91.9
100. 99.7 99.9
85.7
100. 100.
82.2
LGrad
LSUN / ProGAN
72 / 72
100.
91.2
87.3
81.8
82.2
80.6
76.9
66.1
81.1
61.5
68.8
74.1
56.2
57.2
76.1
Ojha et al.
LSUN / ProGAN
360 / 360
100.
93.9
90.9
98.2
96.0
58.4
86.7
80.8
85.7
89.5
92.9
80.6
77.8
85.1
86.9
DIRE-1
LSUN-Bed / ADM
40 / 40
50.6
56.9
49.3
99.9
74.1
44.3
75.7
71.4
68.7
39.4
98.9
99.1
99.6
47.1
69.7
DIRE-2
LSUN-Bed / St.GAN
40 / 40
54.2
52.5
44.3
99.6
76.0
41.0
70.1
70.1
69.3
46.9
97.0
98.2
98.3
42.8
68.6
Ours 1k
COCO / Latent
1 / 1
98.7
89.8
91.9
100.
81.2
89.4
80.3
99.9 94.0
87.4
96.8
98.7
93.9
89.2
92.2
Ours 1k+
COCO / Latent
1 / 1
89.2
79.7
83.2
99.7
74.2
84.0
74.9
99.6
82.0
89.5
97.8
99.2
91.8
90.3
88.2
Ours 10k
COCO / Latent
10 / 10
99.8
92.5
92.7
100.
84.4
88.0
83.2
99.9
96.2
91.8
97.6
99.2
95.2
88.8
93.5
Ours 10k+
COCO / Latent
10 / 10
87.1
81.8
81.9
99.8
73.5
85.7
79.1
99.3
82.5
90.6
98.4
99.7
93.2
90.8
88.8
Table 3: Comparison with SOTA methods in terms of AUC. For our approach we show four variants: trained on 1k real and 1k fake images; 10k real and 10k fake images; trained on 1k and 1k fake images but including compressed/resized images (1k+) and trained on 10k and 10k fake images but including compressed/resized images (10k+). Results obtained on the very same dataset used for training are in light gray, while entries in bold underline the best OOD performance for each dataset. We also report the mean AUC by averaging the AUC scores over all datasets.
GAN family Diffusion family AVG
Pro
Style
Style
Style
Giga
Score
ADM GLIDE
Latent
Stable
DeepFl.
Ediff-I DiT SDXL
Method
real/synth. training
size (k)
GAN
GAN2
GAN3
GANT
GAN
SDE
Diff.
Diff.
IF
Wang et al.
LSUN / ProGAN
360 / 360
100.
69.2
56.2
71.5
52.8
50.4
50.3
50.7
51.3
50.1
50.0
50.4
50.1
49.8
57.3
PatchFor.
CelebA,FF / various
84 / 272
64.8
50.9
50.1
51.1
51.3
50.0
50.3
50.5
51.3
50.2
50.0
50.0
51.2
50.0
51.5
Grag. et al.
LSUN / ProGAN
360 / 360
100. 96.6 92.6
91.9
69.8
83.8
54.5
56.8
67.5
48.9
48.9
52.1
49.2
51.4
68.9
Liu et al.
LSUN / ProGAN
360 / 360
100.
96.2
78.6
89.4
93.3 89.2
67.2
73.2
84.3
65.3
50.5
92.0
64.3
39.6
77.4
Corvi et al.
COCO,LSUN / Latent
180 / 180
50.0
50.7
50.0
52.0
57.7
50.0
54.5
57.3
99.8 99.7
64.3
50.9
99.8 99.7
66.9
LGrad
LSUN / ProGAN
72 / 72
100.
77.2
62.3
68.5
74.7
60.3
61.2
55.4
69.3
61.5
60.9
62.6
53.2
58.7
66.1
Ojha et al.
LSUN / ProGAN
360 / 360
99.9
77.3
82.4
87.1
81.2
57.1
60.9
57.1
68.1
72.6
65.7
66.0
55.4
68.0
71.3
DIRE-1
LSUN-Bed / ADM
40 / 40
51.4
50.8
50.0
96.5
73.1
50.0
73.2
73.2
73.3
49.7
95.3 95.6
94.9
49.9
69.8
DIRE-2
LSUN-Bed / St.GAN
40 / 40
51.4
50.7
50.0
82.0
66.6
50.0
65.9
65.9
66.4
49.6
80.7
81.5
82.2
49.6
63.7
Ours 1k
COCO / Latent
1 / 1
69.7
72.9
82.0
98.9
69.0
76.6
71.2
97.8 86.9
70.9
63.2
80.3
81.0
60.2
77.2
Ours 1k+
COCO / Latent
1 / 1
68.4
70.3
73.5
96.2
62.9
72.8
65.8
94.7
74.5
77.1
81.6
91.9
78.0
75.2
77.4
Ours 10k
COCO / Latent
10 / 10
90.1
83.0
84.0
98.7
66.5
75.6
72.3
93.7
89.9
72.9
59.5
78.8
76.8
54.4
78.3
Ours 10k+
COCO / Latent
10 / 10
73.8
72.3
71.4
95.0
60.8
75.6
68.2
90.5
74.4
77.3
81.7
92.4
77.4
70.0
77.2
Table 4: Comparison with SOTA methods in terms of Accuracy, fixed 0.5 threshold. Results obtained on the very same dataset used for training are in light gray, while entries in bold underline the best OOD performance for each dataset. We also report the mean Accuracy by averaging the Accuracy scores over all datasets.
5 Comparison with the state-of-the-art
In this Section we perform an extensive comparison with SoTA methods in different scenarios. For our approach we show four variants: trained on 1k real and 1k fake images; 10k real and 10k fake images; trained on 1k and 1k fake images but including compressed/resized images (1k+) and trained on 10k and 10k fake images but including compressed/resized images (10k+). In our analysis we only include SoTA techniques with code and/or pre-trained models publicly available on-line in order to ensure a fair comparison. We considered the following approaches:
  1. (a)
    Wang et al. [71] is a CNN detector based on ResNet50 and represents a reference in the research community. This work also introduced the large dataset (LSUN/ProGAN) extensively adopted for model training in subsequent works.
  2. (b)
    Grag et al. [33] proposes a simple modification to the ResNet50 architecture which allows to better preserve low-level forensic traces and is trained on the same dataset introduced in [71].
  3. (c)
    PatchForensics [9] develops a fully-convolutional classifier based on local patches with limited receptive fields over an XceptionNet backbone.
  4. (d)
    Liu et al. [47] is a detector that exploits the inconsistency between real and fake images in the representations of the learned noise patterns. This spatial information is combined with frequency information to improve the classification.
  5. (e)
    LGrad [69] works on the gradients extracted through a pre-trained CNN model in order to filter out the content of the image and transform a data-dependent problem into a transformation-model dependent problem.
  6. (f)
    Corvi et al. [14] performs strong augmentation to gain robustness and increase generalization and is trained on a large dataset of latent diffusion models.
  7. (g)
    Ojha et al. [56] uses CLIP-based features and perfoms detection using the CLIP feature space. Both nearest neighbor and linear probing are tested, even if the latter gives better performance.
  8. (h)
    DIRE [72] is based on the idea that synthetic images can be reconstructed better than real images by a pre-trained model. To this end it is proposed to train a ResNet-50 in two different ways, using ADM images (DIRE 1) and StyleGAN images (DIRE 2), respectively.
Comparison over GAN- and DM-based generators. In Tab. 3 and Tab. 4 we show results in terms of AUC and Accuracy, respectively, over 14 GAN-based and DM-based generators proposed in the scientific literature and perfectly known. For each method, results obtained on the very same dataset used for training are in light gray, to signal an ID scenario, of little interest for our analysis. Entries in bold, instead, underline the best OOD performance for each dataset. Focusing on AUC data, Tab. 3, we observe that methods trained on a GAN dataset work generally well on other dataset of the same family, and much worse on DM datasets. Then, the roles switch for methods trained on DM datasets. This is nothing new, as generators of the two families share architectural details that leave similar traces on the generated images. No SoTA method works uniformly well on all datasets. On the contrary, the proposed lightweight CLIP-based detector provides a uniformly good performance, and the version with 10k+10k reference images without augmentation outperforms the best competitor (notably, again a CLIP-based method) by +6.6%, obtaining an average AUC well above 93.5% a truly impressive accomplishment.
In terms of Accuracy, Tab. 4, performance lowers significantly, and in some cases dramatically. The proposed CLIP-based detectors keep providing the best performance, generally much better than SoTA methods, except for Liu’s method that performs almost at the same level.
Comparison over commercial tools. However, the performance of Liu’s method, as well as that of several other methods, drops catastrophically when considering the Commercial Tools (DALL·E2, DALL·E3, Midjourney, FireFly) both in terms oa AUC and Accuracy, see Tab. 5. Needless to say, this is the most realistic and interesting scenario, often with images of unknown origin and no prior information on the possible generation process. In this situation, most methods provide unreliable decisions. Corvi’s method works surprisingly well on some datasets, maybe generated by diffusion models similar to Latent. However, only the proposed CLIP-based approach provides a stable good performance in all cases. Note that, now, the versions with resized and recompressed images in the reference set perform significantly better than those without augmentation.
Robustness to perturbations. The trend observed over unknown models is further accentuated when images are subject to some post-processing, Tab. 6. These impairments tend to wash out forensic traces, to the point that most SoTA methods become basically useless with coin-tossing performance, especially on the unknown commercial models. Again, the CLIP-based detectors provide a more than decent performance, further supporting our conjecture on the high-level nature of the forensic traces they rely upon.
Commercial tools
DALL·E2
DALL·E3
Midjourney
FireFly
Average
Method
AUC/Acc
AUC/Acc
AUC/Acc
AUC/Acc
AUC/Acc
Wang et al.
64.8 / 50.2
10.9 / 49.4
40.2 / 49.5
84.8 / 50.3
50.2 / 49.8
PatchFor.
41.9 / 50.0
52.7 / 50.0
57.8 / 50.0
49.4 / 50.0
50.5 / 50.0
Grag. et al.
58.3 / 49.2
2.4 / 36.9
43.1 / 40.5
63.5 / 47.5
41.8 / 43.6
Liu et al.
70.4 / 66.4
0.2 / 30.8
40.7 / 43.3
11.8 / 31.0
30.8 / 42.9
Corvi et al.
69.4 / 50.1
60.8 / 50.0
100. / 100.
98.0 / 51.1
82.1 / 62.8
LGrad
58.6 / 55.5
37.9 / 47.8
56.3 / 57.9
40.6 / 47.5
48.3 / 52.1
Ojha et al.
95.2 / 83.5
36.4 / 47.3
66.2 / 52.4
97.5 / 89.9
73.8 / 68.3
DIRE-1
44.7 / 50.0
47.6 / 50.0
51.0 / 49.7
57.4 / 49.9
50.2 / 49.9
DIRE-2
41.0 / 49.9
49.6 / 50.0
47.8 / 49.8
43.0 / 49.8
45.3 / 49.9
Ours 1k
89.1 / 62.2
77.6 / 53.4
74.3 / 53.2
77.1 / 52.1
79.5 / 55.2
Ours 1k+
83.9 / 68.2
94.8 / 83.1
79.7 / 64.7
86.0 / 68.4
86.1 / 71.1
Ours 10k
88.8 / 57.8
82.9 / 51.1
75.4 / 51.2
79.9 / 50.9
81.7 / 52.8
Ours 10k+
84.5 / 63.9
94.4 / 76.6
83.8 / 62.7
87.6 / 64.5
87.6 / 67.0
Table 5: AUC/Accuracy on Commercial Tools. We show AUC and Accuracy on very recent synthetic generators: DALL·E2, DALL·E3, Midjourney v5 and Firefly.
Families of Generators
GAN
Diffusion
Comm. Tools
Average
Method
AUC/Acc
AUC/Acc
AUC/Acc
AUC/Acc
Wang et al.
78.9 / 62.1
59.3 / 50.4
44.1 / 49.9
60.8 / 54.1
PatchFor.
54.3 / 50.0
64.8 / 50.3
62.9 / 50.1
60.7 / 50.1
Grag. et al.
89.2 / 68.5
67.9 / 50.8
47.3 / 50.0
68.1 / 56.4
Liu et al.
54.7 / 51.6
56.3 / 51.1
53.0 / 50.7
54.6 / 51.1
Corvi et al.
66.7 / 54.1
76.1 / 62.1
52.1 / 50.1
65.0 / 55.4
LGrad
51.2 / 51.0
48.0 / 50.5
49.4 / 50.4
49.5 / 50.7
Ojha et al.
86.9 / 74.9
73.7 / 59.8
49.6 / 50.6
70.1 / 61.8
DIRE-1
48.9 / 50.3
51.0 / 50.5
50.1 / 49.9
50.0 / 50.2
DIRE-2
49.1 / 50.0
52.3 / 52.5
45.3 / 49.9
48.9 / 50.8
Ours 1k
76.4 / 68.7
81.2 / 71.4
74.3 / 61.9
77.3 / 67.3
Ours 1k+
75.3 / 67.8
87.8 / 77.6
84.2 / 66.0
82.4 / 70.4
Ours 10k
76.8 / 60.7
82.7 / 59.6
73.8 / 50.5
77.8 / 56.9
Ours 10k+
78.2 / 67.9
88.7 / 77.0
83.0 / 65.6
83.3 / 70.2
Table 6: AUC/Accuracy in the presence of post-processing. All the images have been randomly cropped, resized and compressed to simulate a realistic scenario over the web.
6 Discussion
In this work, we have shown a series of good features of CLIP that can help detecting real from synthetic content. Through an extensive experimental analysis we found that:
  • CLIP features support a much higher generalization ability than what had been discovered so far. By leveraging just a few examples, not even belonging to the generator under test, a simple CLIP-based detector achieves top performance on a large variety of generators and in the most challenging conditions.
  • Maximizing the diversity of reference CLIP features impacts positively on performance, even when a relatively large number of examples is considered.
  • It is well known that CLIP descriptive power descends from its huge pre-training set. Further increasing this set keeps boosting the performance, up to $10\%$.
  • Experimental evidence seems to show that CLIP features, even when tuned to forensic applications, are mostly independent from low-level forensic traces. This opens the way to interesting developments. On one hand, one can gain immunity to possible malicious attacks working on low-level artifacts. On the other hand, there is room for further performance improvement by suitable fusion with classic detectors.
Overall, the present study provides a number of interesting insights but leaves much room for future work. We believe that new and better forensic methods can be proposed on top of CLIP. One major area of study can be to develop few-shot methods that adapt the detector on the fly to the situation of interest. Future work should also consider interpretability, to begin with understanding which forensic features is the detector exploiting to make its decisions.
Acknowledgment.
We gratefully acknowledge the support of this research by a TUM-IAS Hans Fischer Senior Fellowship, a TUM-IAS Rudolf Mößbauer Fellowship and a Google Gift. This material is also based on research sponsored by the Defense Advanced Research Projects Agency (DARPA) and the Air Force Research Laboratory (AFRL) under agreement number FA8750-20-2-1004. The U.S. Government is authorized to reproduce and distribute reprints for Governmental purposes notwithstanding any copyright notation thereon. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either expressed or implied, of DARPA or the U.S. Government. In addition, this work has received funding by the European Union under the Horizon Europe vera.ai project, Grant Agreement number 101070093. It is also supported by the PREMIER project, funded by the Italian Ministry of Education, University, and Research within the PRIN 2017 program. Finally, we want to thank Koki Nagano for useful discussions, and Tero Karras, Yogesh Balaji, Ming-Yu Liu for sharing data for StyleGAN-T and eDiff-I experiments.
Supplemental Material
In this appendix, we report additional ablation studies (Sec. A) and robustness analysis (Sec. B). We also enlarge our initial dataset with additional synthetic generators and carry out further experiments on generalization (Sec. C). Furthermore, we conduct more experiments on the ability of our approach and SoTA methods to deal with possible attacks from malicious users (Sec. D). Finally, we show some experiments carried out on social networks in a few-shot scenario (Sec. E).
Figure 6: Average AUC for our approach considering different classifiers and different features on the original dataset. Top: 1k images, Bottom: 10k images.
Figure 7: Average AUC for our approach considering different classifiers and different features on the dataset after random compression and resizing. Top: 1k images, Bottom: 10k images.
Appendix A Additional ablation study
In Section 4 of the paper we describe the CLIP-based detection procedure used in all experiments. Two important design choices outlined in the paper are that i) we extract features from the next to last layer of the CLIP ViT L/14 architecture; and ii) we use a SVM classifier, designed based on a limited set of reference features. Here we test some alternative solutions, that is:
  • extracting features from the last layer of the architecture;
  • using other classifiers: Logistic Regression (LR), Mahalanobis distance (MAH), Gaussian Naive Bayes classifier (GNB), Soft voting k-Nearest Neighbor (SNN) [54].
In Figures 6 and 7, we show the results in terms of average AUC over the dataset described in Section 3 (main paper) using both original images and images impaired by common post-processed steps such as recompression and resizing. For all detectors, we consider the versions 1k and 10k (1000 and 10000 reference images, respectively).
In all cases, the SVM classifier seems to ensure the best performance, followed closely by the Logistic Regression, while the other classifiers provide a less consistent performance. Using features from the next-to-last layer is almost always preferable to using features from the last layer, and this happens always with the SVM classifier. These results motivated our design choices.
Figure 8: Robustness analysis in terms of AUC carried out on the Stable Diff. 2, SDXL, DALL·E 3, DALL·E 2, Midjourney, and Firefly generation models of the SynthBuster dataset [4].
Appendix B Additional robustness analysis
In Section 5 of the paper, in Tab. 6, we provide some results on the robustness of SoTA and proposed methods in the presence of image impairments. Due to lack of space, however, we show only some very synthetic results, averaged on generic post-processed images. Here, we analyze robustness in more detail as a function of the JPEG Quality Factor, which ranges from 100 to 60, and the resizing scale, which ranges from 125% to 25%. In addition, we consider also another compression method, WEBP, which is gaining popularity on social networks and has been rarely been used for augmentation. Images are from the SynthBuster dataset [4], generated by Stable Diffusion 2, SDXL, DALL·E 2, DALL·E 3, Midjourney, and Firefly. We show results (AUC) only for the most competitive (in terms of robustness) SoTA methods, Corvi and Ojha. For our method we consider again four variants similarly to the main paper: the versions using 1k or 10k real and fake reference images, and the corresponding versions, 1k+ and 10k+, where data have been augmented with randomly compressed/resized images, while keeping constant the total number of images.
Ojha suffers some performance losses, up to 10 points, in the presence of compression, be it JPEG or WEBP, and a much more rapid decline with resizing, to the point of becoming useless in the presence of a 25% rescaling. Corvi was trained with strong augmentation and, in fact, is not affected at all by JPEG compression, while it presents acceptable losses with strong WEBP compression (not seen in training), and strong resizing. The most remarkable outcome of this experiment, however, is the impressive robustness of the CLIP-based detectors. Both the 1k+ and 10k+ versions, the ones with augmentation, are basically insensitive to compression, no matter JPEG or WEBP, and resizing. The versions without augmentation suffer some loss of performance but not nearly as dramatic as for the reference methods, and remain effective in all conditions.
Generator
modality
LSUN [76]
FFHQ [41]
ImageNet [18]
COCO [46]
Resolution
BigGAN [7] c $256^{2}$, $512^{2}$
EG3D [10] u $512^{2}$
Diff.ProjectedGAN [73] u $256^{2}$
GALIP [70] t $224^{2}$
Taming Transf. [27] u,c $256^{2}$
DALL·E mini [17] t $256^{2}$
DDPM [34] u $256^{2}$
Deepfloyd-IF II stage [44] t $256^{2}$
Table 7: Image generators used in our experiments: GAN-based, VQ-GAN-based and DM-based. The generation modalities, unconditional (u), conditional (c), and text-to-image (t), is reported in the second column. The last column reports the resolution of the images in the dataset.
Appendix C Further generalization results
In this Section we extend our generalization analysis to additional data. In the main paper we included 6,000 real images and 18,$0$ fake images from 18 models, now we consider 8 more generators with 8,000 additional synthetic images for our test set (Tab. 7). The StyleGAN-T [63], eDiff-I [3], and GigaGAN [39] images are provided by the authors of the original papers. Instead, we ourselves generated images for Score-SDE [68], Stable Diff. [62], DiT [57], Deepfloyd-IF [44], Diff.ProjectedGAN [73], GALIP [70], and DDPM [34] using the pre-trained models available on-line. The other generated images come from three public datasets [71, 14, 4]. To avoid possible biases, performance metrics are evaluated using for each fake dataset images drawn from the real dataset used in training, indicated in the central columns of Table 1 of main paper and Table 7 of this supplemental.
AUC/Acc BigGAN EG3D
Diff. Proj.
GALIP
Taming
DALL·E
DDPM
Deepfloyd-IF
AVG
Method
GAN
Transf.
Mini
II stage
[71] Wang et al.
92.7 / 66.1
94.4 / 59.2
89.8 / 52.1
89.7 / 57.4
54.3 / 51.7
62.5 / 51.8
31.6 / 50.1
43.1 / 50.1
69.8 / 54.8
[9] PatchFor.
85.5 / 52.5
69.8 / 50.0
92.6 / 61.7
98.2 / 73.4
71.2 / 51.0
83.8 / 51.4
98.4 / 50.2
83.4 / 50.0
85.4 / 55.0
[33] Grag. et al.
98.7 / 94.2
98.9 / 93.8
100. / 72.5
96.2 / 79.5
90.3 / 76.4
83.4 / 62.5
49.7 / 45.1
71.2 / 61.3
86.0 / 73.2
[47] Liu et al.
94.7 / 81.3
99.0 / 86.3
99.5 / 84.9
94.3 / 78.4
95.4 / 78.9
98.4 / 88.1
22.8 / 49.6
97.4 / 86.8
87.7 / 79.3
[14] Corvi et al.
83.4 / 51.8
25.2 / 50.0
96.6 / 71.2
87.7 / 50.9
99.3 / 89.5
99.7 / 83.8
100. / 90.3
68.5 / 50.8
82.5 / 67.3
[69] LGrad
77.2 / 69.0
68.8 / 59.8
99.5 / 90.1
56.7 / 55.1
74.1 / 64.5
67.3 / 59.9
9.8 / 16.4
75.0 / 62.5
66.1 / 59.7
[56] Ojha et al.
99.6 / 96.4
92.6 / 82.5
97.4 / 75.2
98.6 / 89.9
94.1 / 84.8
97.1 / 84.9
77.7 / 68.2
60.8 / 50.0
89.7 / 79.0
[72] DIRE-1
99.8 / 95.3
50.1 / 50.0
49.8 / 51.6
100. / 96.7
73.1 / 72.4
99.7 / 96.5
23.1 / 50.0
99.4 / 95.8
74.4 / 76.0
[72] DIRE-2
98.6 / 82.4
46.1 / 50.0
50.2 / 51.4
99.3 / 81.8
77.0 / 66.2
98.7 / 81.8
14.0 / 49.8
95.6 / 80.5
72.4 / 68.0
Ours 1k
96.7 / 85.2
86.3 / 73.8
99.3 / 69.7
100. / 99.7
99.3 / 94.3
100. / 98.9
94.7 / 86.0
99.7 / 90.6
97.0 / 87.3
Ours 1k+
90.9 / 77.4
75.5 / 64.7
89.4 / 68.4
99.9 / 98.2
93.4 / 82.3
99.6 / 94.0
94.6 / 87.7
99.7 / 95.1
92.9 / 83.5
Ours 10k
98.2 / 85.9
90.8 / 81.4
99.8 / 90.1
100. / 99.9
99.6 / 94.3
100. / 99.1
94.9 / 88.3
99.8 / 88.1
97.9 / 90.9
Ours 10k+
91.7 / 75.0
80.7 / 72.5
88.9 / 74.6
100. / 97.8
92.5 / 77.2
99.8 / 93.5
95.9 / 79.5
99.8 / 92.8
93.7 / 82.8
Table 8: Comparison with SOTA methods on additional data. The results are in terms of AUC/Accuracy. The entries in bold underline the best performance for each dataset. For our approach we show four variants: trained on 1k real and 1k fake images; 10k real and 10k fake images; trained on 1k and 1k fake images but including compressed/resized images (1k+) and trained on 10k and 10k fake images but including compressed/resized images (10k+).
AUC/Acc BigGAN EG3D
Diff. Proj.
GALIP
Taming
DALL·E
DDPM
Deepfloyd-IF
AVG
Method
GAN
Transf.
Mini
II stage
[71] Wang et al.
82.5 / 55.5
84.7 / 52.2
80.8 / 53.2
92.2 / 59.2
66.2 / 50.6
66.7 / 50.5
69.6 / 50.2
47.9 / 49.9
73.8 / 52.7
[9] PatchFor.
58.7 / 50.3
60.3 / 50.0
55.6 / 50.0
72.3 / 50.3
50.4 / 50.1
71.0 / 50.1
82.9 / 51.0
69.7 / 50.0
65.1 / 50.2
[33] Grag. et al.
97.6 / 74.4
92.7 / 56.5
99.2 / 65.6
99.1 / 79.8
83.6 / 53.3
89.6 / 54.0
74.5 / 49.4
54.3 / 50.0
86.3 / 60.4
[47] Liu et al.
57.7 / 52.0
51.6 / 50.4
62.9 / 54.1
65.2 / 52.3
49.2 / 50.2
49.1 / 50.7
58.3 / 49.1
64.4 / 52.0
57.3 / 51.4
[14] Corvi et al.
77.3 / 53.0
64.5 / 59.7
93.3 / 59.8
87.4 / 51.5
94.4 / 76.7
97.4 / 74.8
74.7 / 59.6
70.4 / 50.9
82.4 / 60.7
[69] LGrad
53.5 / 51.6
53.2 / 51.1
56.9 / 51.1
57.0 / 50.3
51.3 / 51.1
41.7 / 49.1
56.0 / 49.7
38.7 / 48.8
51.0 / 50.4
[56] Ojha et al.
94.8 / 84.9
79.3 / 61.5
84.7 / 68.5
95.7 / 84.5
92.9 / 81.8
89.1 / 72.8
90.8 / 79.1
59.1 / 49.5
85.8 / 72.8
[72] DIRE-1
47.2 / 49.5
35.2 / 49.3
35.2 / 49.6
47.4 / 50.6
47.3 / 49.7
54.6 / 51.1
38.5 / 49.4
64.6 / 53.2
46.2 / 50.3
[72] DIRE-2
44.6 / 47.0
33.6 / 47.1
37.5 / 44.9
46.2 / 47.2
43.3 / 46.8
53.1 / 50.0
42.2 / 47.5
66.3 / 59.3
45.9 / 48.7
Ours 1k
84.0 / 74.8
64.4 / 59.5
75.3 / 68.3
98.5 / 93.0
85.3 / 75.0
96.5 / 89.0
76.7 / 66.6
90.6 / 80.1
83.9 / 75.8
Ours 1k+
88.9 / 78.2
60.0 / 54.6
70.5 / 60.9
99.8 / 98.3
86.5 / 77.3
98.9 / 93.6
96.1 / 84.7
99.4 / 96.2
87.5 / 80.5
Ours 10k
84.4 / 59.1
70.9 / 63.3
68.8 / 55.6
99.2 / 79.7
85.7 / 67.7
97.5 / 68.7
64.6 / 57.0
95.1 / 62.2
83.3 / 64.1
Ours 10k+
91.3 / 79.4
68.7 / 59.7
68.7 / 61.7
99.9 / 99.0
86.2 / 71.2
99.4 / 93.6
92.7 / 62.4
99.8 / 97.8
88.3 / 78.1
Table 9: Comparison with SOTA methods on additional data in the presence of post-processing. The results are in terms of AUC/Accuracy. The entries in bold underline the best performance for each dataset.
In Tab. 8 we show the results in terms of AUC and Accuracy on such data for the four version of the proposed CLIP-based detector and for the SoTA methods described in the main paper (Section 5): Wang et al. [71], Grag et al. [33], PatchForensics [9], Liu et al. [47], LGrad [69], Corvi et al. [14], Ojha et al. [56], DIRE [72].
[…]
Appendix D Attack scenario
In this section, we conduct some additional experiments on a specific attack scenario, where a malicious user takes real images and passes them through the autoencoder of a specific generative model. The objective is to insert the low-level traces typical of synthetic architectures in real images in order to fool forensic detectors and make everything appear as synthetic. To this end, we took 1000 real images from the RAISE dataset [16] and passed them through the autoencoders of three different versions of diffusion models: Stable Diffusion 1.4, Stable Diffusion 2.0 [62] and Stable Diffusion XL [58]. In Fig. 9 we show that such process insert exactly the same forensic clues in the real images and they appear indistinguishable from the generated ones, without altering their visual appearance (see Fig. 5 of the main paper).
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 9: Top (from left to right): Fourier spectra of the noise residuals for synthetic images generated with Stable Diffusion 1.4, Stable Diffusion 2.0 and Stable Diffusion XL. Bottom: Fourier spectra of the noise residuals for the real images that passed through the corresponding autoencoder of the synthetic generators. We can observe that this operation has inserted in the real images exactly the same artifacts present in the synthetic ones.
In Table 10 we report results in terms of accuracy before and after the attack for our method and the best competitors. Before the attack, the detector must discriminate between real and generated images while, after the attack, it must discriminate between laundered real images and generated ones. We notice that before the attack the method proposed by Corvi et al. can achieve almost perfect results. However, it is very sensitive to the insertion of low-level traces and provides totally wrong results on the attacked data. For Ojha et al the performance also worsen, with an average 7% loss. Instead, for our method the Accuracy is almost unaffected by the attack, with an average loss below 1.3%.
Based on these results, it appears that the proposed CLIP-based detector does not rely on the same low-level traces exploited by most of the current detectors. This ensures higher robustness to common attacks and opens the way to the use of suitable fusion strategies. However, there is much room for further analysis, to begin with identifying which traces our detector is actually exploiting.
Accuracy before the attack after the attack
Method
St.Diff.
St.Diff.
SDXL
St.Diff.
St.Diff.
SDXL
1.4
2.0
1.4
2.0
Corvi et al.
100.
99.2
99.7
50.0
49.2
53.7
Ojha et al.
69.9
75.6
68.0
59.0
69.0
64.5
Ours 1k
63.6
64.3
60.2
63.6
63.8
59.2
Ours 1k+
76.8
77.0
75.2
77.0
75.5
72.6
Ours 10k
65.8
60.5
54.4
65.9
60.2
54.0
Ours 10k+
78.8
72.8
70.0
79.1
71.5
68.7
Table 10: Results in terms of Accuracy before and after the laundering attack on real images. Before the attack all methods achieve a very good performance. After the attack the performance of our method is almost unchanged, while significant losses are observed for the best reference methods, especially for Corvi et al. that goes from perfect detection to random guess.
Appendix E Few-shot analysis in the wild
In this section we present an experiment in a few-shot scenario. The idea is to explore the ability of our detector to work with very limited data and adapt to a situation where only a few real and synthetic examples are available. We downloaded real and synthetic images from Twitter and considered three different generators DALL·E 3, Midjourney and Firefly (some examples can be seen in Fig. 10). To understand which generator was used to create a specific image, we relied on tags and annotations present on the social network.
In our few-shot scenario, we take $10$ real images and $10$ generated images as examples from a specific model and test on all the others images (note that we present average results on 1000 runs). Results are reported in Tab. 11 in terms of AUC and Accuracy. The availability of just 10+10 images of the target data provides an impressive performance boost with respect to the same method trained on a larger dataset, but not aligned with the test data (real images come from COCO, synthetic images from Latent diffusion models). It is important to underline that this is a realistic scenario, in which one is called to decide on images generated by a new method and a few sample images are available as a support. We believe that this can represent an interesting direction for the application in the wild or to easily adapt a detector to more challenging situations where some prior information is available.
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 10: Some examples of images downloaded from the social network Twitter. From left to right: a real image, synthetic images from DALL·E 3 [6], Firefly [30], Midjourney [52].
AUC/Acc Twitter
Method
DALL·E3
Midjourney
FireFly
Average
Ours 1k
71.9 / 51.1
70.3 / 54.0
77.4 / 52.4
73.2 / 52.5
Ours 1k+
83.6 / 72.3
75.5 / 67.3
89.6 / 80.5
82.9 / 73.4
Ours 10k
68.2 / 50.5
72.0 / 51.5
70.7 / 51.5
70.3 / 51.2
Ours 10k+
67.0 / 60.0
78.3 / 71.0
74.3 / 68.5
73.2 / 66.5
Ours Few-Shot
97.7 / 90.6
91.9 / 82.5
96.6 / 88.4
95.4 / 87.2
Table 11: We show results in a few-shot experiment, where we suppose to know in advance 10+10 real/fake images from a specific generator and compare them with our basic approach trained on a larger set of real/fake images from misaligned data. As can be seen, even with just 10 images it is possible to adapt the model and obtain a large improvement.

### v2

@@
 Abstract
-Aim of this work is to explore the potential of pre-trained vision-language models (VLMs) for universal detection of AI-generated images.
+The aim of this work is to explore the potential of pre-trained vision-language models (VLMs) for universal detection of AI-generated images.
 We develop a lightweight detection strategy based on CLIP features and study its performance in a wide variety of challenging scenarios.
-We find that, unlike previous belief, it is neither necessary nor convenient to use a large domain-specific dataset for training.
-On the contrary, by using only a handful of example images from a single generative model, a CLIP-based detector exhibits a surprising generalization ability and high robustness across several different architectures, including recent commercial tools such as Dalle-3, Midjourney v5, and Firefly.
-We match the SoTA on in-distribution data, and improve largely above it in terms of generalization to out-of-distribution data (+6% in terms of AUC) and robustness to impaired/laundered data (+13%).
+We find that, contrary to previous beliefs, it is neither necessary nor convenient to use a large domain-specific dataset for training.
+On the contrary, by using only a handful of example images from a single generative model, a CLIP-based detector exhibits surprising generalization ability and high robustness across different architectures, including recent commercial tools such as Dalle-3, Midjourney v5, and Firefly.
+We match the state-of-the-art (SoTA) on in-distribution data and significantly improve upon it in terms of generalization to out-of-distribution data (+6% AUC) and robustness to impaired/laundered data (+13%).
 Our project is available at https://grip-unina.github.io/ClipBased-SyntheticImageDetection/
 1 Introduction
-Synthetic images have by now left the research laboratories and are flooding the real world.
+Synthetic images have by now left research laboratories and are flooding the real world.
 The latest versions of popular image editing tools, such as Adobe Photoshop and Microsoft Paint, come with easy-to-use AI-powered generative tools that allow even novice users to edit and generate visual content at will.
-Thanks to diffusion-based generative models, it is not only the quality of the images generated that has improved enormously, but also the flexibility in using these tools.
-By issuing a few textual commands, it is possible to easily obtain the desired image.
-While this represents a great opportunity for visual art applications, it is also the paradise of disinformation professionals who can design their attacks with unprecedented power and flexibility [8, 5, 25].
-And, of course, a nightmare for all those trying to combat the spread of fake news and orchestrated disinformation campaigns: journalists, fact-checkers, law enforcement agencies, governments.
-Therefore, there is high demand for automatic tools that help establishing the authenticity of a media asset.
+Thanks to diffusion-based generative models, it is not only the quality of the generated images that is greatly improved, but also the flexibility in using these tools.
+By issuing a few text commands you can easily obtain the desired image.
+While this represents a great opportunity for visual arts applications, it is also the paradise of disinformation professionals who can design their attacks with unprecedented power and flexibility [8, 5, 26].
+And, of course, a nightmare for all those trying to combat the spread of fake news and orchestrated disinformation campaigns: journalists, fact-checkers, law enforcement, governments.
+Therefore, there is a high demand for automatic tools that help establish the authenticity of a media asset.
 Figure 1: Area Under ROC Curve (AUC %) on unseen synthetic generators ($x$-axis) and on post-processed data ($y$-axis).
-The first number measures the generalization ability of the detector, the second its robustness to possible impairments.
-Circle area proportional to training set size.
-The performance is measured over 18 different synthetic models.
+The first number measures the generalization ability of the detector, the second measures its robustness to possible impairments.
+Circle area is proportional to training set size.
+Performance is measured over 18 different synthetic models.
 Our CLIP-based detector largely outperforms all SoTA methods with very limited training data.
 It is well known that each synthesis model leaves its own peculiar traces in all generated images, subtle traces that give rise to so called artificial fingerprints, that can be exploited for forensic analyses [50, 79].
-This was first observed for methods based on generative adversarial networks (GAN) [79, 22] but holds with obvious differences for all generation approaches, including the more recent diffusion models (DM) [13].
+This phenomenon was first observed for methods based on generative adversarial networks (GAN) [81, 23] but holds, with obvious differences, for all generation approaches, including the more recent diffusion models (DM) [13].
 One of the main challenges for current forensic detectors is the ability to generalize to new and unseen generative methods.
-Indeed, the in-distribution (ID) scenario, with perfectly aligned training and test set, is rarely met in practice.
-Test images are often generated by new approaches, unseen architectures, or even known architectures re-trained under different conditions, all different flavors of the out-of-distribution (OOD) scenario.
-In addition, most of the times, images are downloaded from some social networks, where they undergo transformations such as compression and resizing (even multiple times), processes that tend to wash out the tiny traces so precious for detection.
+Indeed, the in-distribution (ID) scenario, with perfectly aligned training and test sets, is rarely met in practice.
+Test images are often generated by new approaches, unseen architectures, or even known architectures re-trained under different conditions, all representing different flavors of the out-of-distribution (OOD) scenario.
+Additionally, most images are downloaded from social networks, where they undergo transformations such as compression and resizing (sometimes multiple times), processes that tend to wash out the tiny traces so precious for detection.
 In [73], it is shown that training diversity and intense augmentation are crucial for generalization.
-A ResNet-50 detector is trained on a single but very diverse dataset of about 360k ProGAN images [40].
+A ResNet-50 detector is trained on a single but highly diverse dataset of about 360k ProGAN [41] images.
 Results on OOD images generated by other GAN-based generators turned out to be surprisingly good.
-The authors themselves, however, suggest that structural similarities between GAN-based generators are the likely reason for such good performance.
+However, the authors themselves attribute this performance to the structural similarities between GAN-based generators.
 Indeed, results are not equally good on images generated by recent diffusion-based methods, which present somewhat different generation artifacts [14].
-Generalizing to OOD data is a major issue in deep learning and has been the object of intense research in latest years.
+Generalizing to OOD data is a major issue in deep learning and has been the object of intense research in recent years.
 In this context, the advent of large pre-trained vision-language models has brought about a number of new solutions and exciting results.
@@
 Recently, there have been attempts to exploit the power of VLMs to detect synthetic images [67, 2, 57].
-For example, in [56] CLIP (Contrastive Language-Image Pre-Training) [59] is used as a pre-trained feature extractor, and a classifier is then trained on the 360k-image ProGAN dataset proposed in [71], with the same augmentation strategy.
-With respect to [71] based on Resnet50, the performance improves significantly, especially on images generated by diffusion models, showing a very good generalization ability.
-In this work we explore in depth the potential of CLIP for image forensics, propose new decision strategies, and carry out an extensive experimental analysis in challenging real-world scenarios involving a large number of generative models.
+For example, [57] considers the same dataset, augmentation strategy and experimental protocol as in [73] but uses a pre-trained VLM, the Contrastive Language-Image Pre-Training (CLIP) [60], as a feature extractor rather than training a ResNet-50.
+Only the classifier is then learned on the task-specific dataset.
+Compared to [73] the performance improves significantly, especially on images generated by diffusion models, showing excellent generalization ability.
+In this work, we explore in depth the potential of CLIP for image forensics and conduct an extensive experimental analysis in challenging real-world scenarios involving a large number of generative models.
 We find that CLIP-based methods show impressive generalization ability and very good robustness, resulting in large performance improvements compared to the state of the art.
-To obtain such results we avoid any intensive training on domain-specific data, which may introduce unwanted biases and undermine the descriptive power of CLIP features.
+To achieve such results, we avoid any intensive training on domain-specific data, which could introduce unwanted biases and undermine the descriptive power of CLIP features.
 Instead, we rely on a small number of paired real/fake images with the same textual description and use their CLIP features to model the decision space.
-Synthetic images come from a single generator but work well with all other tested data sources.
-The top performance is obtained with just 1000 to 10000 paired images.
-Moreover, only minor decays are observed when this number reduces to 100 or even 10 (see Fig.
+Synthetic images come from a single generator, but results are equally good across all different data sources.
+The top performance is achieved with just 1,000 to 10,000 paired images.
+Moreover, only minor performance decays are observed when this number reduces to 100 or even 10 (see Fig.
 1).
@@
   • We show that CLIP features achieve excellent generalization: by exploiting only a handful of examples, not even belonging to the generator under test, the performances are comparable to those of intensively trained solutions.
-  • We carry out a large set of experiments on diverse synthetic generators and in very challenging conditions and are able to achieve the best performance (on average).
-Our experiments make clear that the features extracted are at least partially orthogonal to those explored by previous methods mostly oriented to extract low-level features.
+  • We carry out a large set of experiments on diverse synthetic generators and in very challenging conditions achieving the best performance on average.
+Our experiments make clear that the features extracted are partially orthogonal to the low-level features used by previous methods.
 2 Related work
-There is an extensive literature on synthetic image detection, mostly concerning images created by generative adversarial networks (GANs) and, more recently, by diffusion models (DMs).
-Some methods look for visible errors, such as asymmetries in faces, wrong perspective, odd shadows [51, 28, 29].
-However, with the rapid progress in image synthesis technology, this kind of problems are solved quickly and appear more and more rarely in modern generation methods.
-Hence, we will focus on methods that exploit “invisible” forensic traces, working either in the spatial or the frequency domain.
-Then we will briefly review the most recent approaches that exploit multimodal features.
+There is an extensive literature on synthetic image detection, primarily focusing on images created by generative adversarial networks (GANs) and, more recently, by diffusion models (DMs).
+Some methods look for visible errors, such as asymmetries in faces, incorrect perspectives, or unusual shadows [52, 29, 30].
+However, with the rapid progress in image synthesis technology, these types of problems are solved quickly and appear less frequently in modern generation methods.
+Therefore, we will focus on methods that exploit “invisible” forensic traces, working either in the spatial or frequency domain.
+Subsequently, we will provide a brief overview of the most recent approaches that exploit multimodal features.
 Spatial domain methods.
-Despite their high visual quality, synthetic images carry with them distinctive traces of the generation process that enable detection and even model identification [49, 77].
-In fact, each generation model inserts a sort of digital fingerprint into all the images created, which depends on its architectural and training details.
-This fingerprint can be easily estimated, Given an adequate number of images generated by the target model, typically a few hundred, it is sufficient to average their noise residuals, extracted through simple denoisers [49] or more sophisticated deep learning-based methods [67, 47].
+Despite their high visual quality, synthetic images carry with them distinctive traces of the generation process that enable detection and even pattern identification [50, 79].
+Each generation model, in fact, inserts a sort of digital fingerprint into all the images it creates, which depends on its architectural and training details.
+This fingerprint can be easily estimated.
+Given a few hundred images generated by the target model, it is sufficient to average their noise residuals, extracted using simple denoisers [50] or more sophisticated methods based on deep learning [68, 48].
 Detectors based on digital fingerprints can be regarded as “few-shot”, as they can deal with new models based on just a few example images.
 Other few-shot solutions have been proposed lately in this field [15, 51, 22, 37].
-However, they all require some prior knowledge of the target models, even if limited to a few images, so they fit a strictly in-distribution (ID) scenario.
-In contrast, our work aims to generalize to new and previously unseen models, and therefore fits into an out-of-distribution (OOD) scenario.
+However, they all require information on the target models, even if limited to a few images, so they fit a strictly in-distribution scenario.
+In contrast, our work aims to generalize to new and previously unseen models, and therefore fits into an out-of-distribution scenario.
 Good generalization is a key requirement of synthetic image detectors.
 Towards this end, it is important to increase diversity in training, which can be pursued by suitable augmentation, by including a large number of different categories [73], or even by model ensembling [49].
-Other studies suggest to work on local patches [9] or combine global spatial information with local features [38].
-In [33] it is shown that to preserve the subtle high-frequency forensic traces, one should avoid any downsampling in the first layers of the network.
-With the same aim, [69] proposes to work on noise residuals rather than the original images, by extracting the gradients with a pre-trained CNN.
-In [12] a systematic study is carried out on transferable forensic features that allows easier generalization.
+Other studies suggest working on local patches [9] or combine global spatial information with local features [39].
+In [34], it is shown that to preserve the subtle high-frequency forensic traces, one should avoid any downsampling in the first layers of the network.
+With the same aim, [70] proposes to work on noise residuals rather than the original images, extracting the gradients with a pre-trained CNN.
+In [12], a systematic study is carried out on transferable forensic features that allows easier generalization.
 The above investigations and proposals, however, only consider GAN-based generators.
@@
 On the other hand, [25] shows that it is possible to achieve decent performance by continuously re-training the detector on images from new generators, as long as the latter share some architectural components with the known ones.
-In [72], inspired to previous work on GAN images [1], a detector specifically tailored to DM images is proposed.
+In [74], inspired by previous work on GAN images [1], a detector specifically tailored to DM images is proposed.
 Images are projected in a latent space and reconstructed with known models to study their features.
-While promising, the approach refers to specific architectures, hence an ID scenario, and has been tested on a limited set of diffusion generative models.
-In the present work, instead, we do not make any assumptions on the test data and include a large variety of both GAN and DM generators including OOD images, and realistic high-resolution images generated by recent diffusion models.
+Although promising, the approach refers to specific architectures, hence an ID scenario, and has been tested on a limited set of generative diffusion models.
+In this work instead, we do not make any assumptions and include a large variety of both GAN and DM generators.
 Frequency domain methods.
@@
 In fact, they are caused by the up-sampling operations used in the generator, which give rise to regular spatial patterns and strong peaks in the Fourier domain.
-Interestingly, the awareness of such weakness has prompted the design of new architectures, like StyleGAN3 [43], which explicitly avoid aliasing and reduce the above mentioned peaks.
+Interestingly, the awareness of such weakness has prompted the design of new architectures, like StyleGAN3 [44], which explicitly avoid aliasing and reduce the above-mentioned peaks.
 In addition to such obvious artifacts, the spectral content of GAN images and real images is also known to differ significantly in the medium and high frequencies.
 Likewise, clear differences in the radial and angular spectral distributions of the DM and real images were observed [13, 77].
-Furthermore, it is worth observing that DM images also may present spectral peaks (see Fig.
+Furthermore, it is worth observing that DM images may also present spectral peaks (see Fig.
 5).
 Of course, frequency domain traces can be used to train simple detectors for ID images.
-Even more interestingly, some authors [79, 37] replicate and modify the synthesis process of known generators by introducing a series of small architectural changes in order to learn to manage also OOD test images.
-A major problem with spectral traces is that they can be easily weakened by laundering operations, both innocent (e.g., image resizing) and malicious counterforensic methods [20, 11].
-Likewise, attackers can easily embed artificial spectral traces in a real image to make it appear as synthetic.
-This is a more general problem of low-level traces, which can be attacked in both ways and through a number of ingenious methods.
-We will show that our approach is indeed immune to such attacks and can handle a variety of distribution shift.
-Exploiting multimodal features.
-The introduction of large language models has elicited intense work in the vision community on trying to exploit large models trained using both images and text [80, 78].
-In image forensics, however, traditionally oriented to the analysis of low-level features, there has been very little work in this direction.
-Only a few papers [66, 2, 56] have tried exploiting pre-trained visual models for the detection of AI-generated images.
-All of them rely on variants of CLIP, however only in [56] it is pursued generalization to OOD data, by nearest neighbor and linear probing in the CLIP feature space.
+Even more interestingly, some authors [81, 38] replicate and modify the synthesis process of known generators by introducing a series of small architectural changes to learn to handle even OOD test images.
+A major problem with spectral traces is their low resilience to laundering operations, both unintentional (e.g.
+image resizing) and intentional (e.g., counterforensic methods [21, 11, 16]).
+Methods based on multimodal features.
+The introduction of large language models has sparked intense work in the vision community attempting to leverage large models trained using both images and text [82, 80].
+In image forensics, however, traditionally oriented towards the analysis of low-level features, there has been very little work in this direction.
+Only a few papers [67, 2, 57] have tried to leverage pre-trained visual models for AI-generated image detection.
+All these studies rely on variants of CLIP.
+However, only in [57] is generalization to OOD data pursued, via nearest neighbor and linear probing in CLIP feature space.
 To this end, a large dataset of fake and real images is used to train the classifier.
-In this work we make a further step in this direction and show that better performance can be achieved even using much less data.
+In this work we take a further step in this direction and show that superior performance can be achieved even using much less data.
 Refer to caption Refer to caption Refer to caption
@@
 3 Datasets and metrics
-The main objective of this work is to study the generalization ability of CLIP-based and conventional methods, that is, assess how the detectors behave on data never seen before, of different provenance, and possibly impaired by various forms of post-processing.
-Therefore, we consider a very large test dataset and try to avoid any possible biases in the experimental validation.
+The main objective of this work is to study the generalization capacity of conventional and CLIP-based methods, i.e.
+to evaluate how detectors behave on data never seen before, of different origins and possibly impaired by various forms of post-processing.
+Therefore, we consider a very large test dataset and take measures to mitigate any potential biases in the experimental validation process.
 Tab.
-1 lists all the generation models used in testing.
-They are grouped in three families: GAN-based (ProGAN, StyleGAN2, StyleGAN3, StyleGAN-T, GigaGAN), DM-based (Score-SDE, ADM, GLIDE, eDiff-I, Latent and Stable Diffusion, DiT, DeepFloyd-IF, Stable Diffusion XL) and commercial tools (DALL·E 2, DALL·E 3, Midjourney v5 and Adobe Firefly), which includes some recent text-driven methods available on-line, whose architecture is not always disclosed.
-These models cover the different generation modalities commonly adopted in this field: unconditional (u), conditional (c), and text-to-image (t), reported in second column of the table.
+1 lists all generation models used during test.
+They are grouped into three families: GAN-based (ProGAN, StyleGAN2, StyleGAN3, StyleGAN-T, GigaGAN), DM-based (Score-SDE, ADM, GLIDE, eDiff-I, Latent and Stable Diffusion, DiT, DeepFloyd-IF, Stable Diffusion XL) and commercial tools (DALL·E 2, DALL·E 3, Midjourney v5 and Adobe Firefly), which include some recent text-based methods available online, whose architecture is not always disclosed.
+The second column of the table specifies the generation modalities employed by these models: unconditional (u), conditional (c), and text-to-image (t).
 The last column reports the resolution of the images in the dataset.
-The central columns, instead, indicate which datasets of real images have been considered to carry out the detection experiments.
+The central columns, indicate which datasets of real images were considered to carry out the detection experiments.
 These have been carefully selected to avoid all foreseeable biases.
-For example, when testing on synthetic images generated by ProGAN, we use as real counterparts images drawn from the LSUN dataset on which the ProGAN model was trained.
-Therefore, if LSUN images shared all the same defect (say, a red ”LSUN” label in a corner) the ProGAN images should spot the same defect, which could not be used to spoil detection performance.
-This hilarious example should not hide the importance of the problem, as biases are way too common in experiments and not always so easy to avoid.
-In other cases, for example when text-driven commercial tools with undisclosed training details are used, we relied on a dataset recently proposed in [4].
-Pristine and synthetic images with similar semantic content are built by extracting textual descriptions from a real dataset (RAISE [16]) and then using them as prompt for generating synthetic ones.
-Overall, we selected $32,000$ real and fake images for testing, which will form the dataset used in all forthcoming experiments.
-To consider a more realistic scenario, we also consider datasets with post-processing operations that simulate images shared over social networks.
-In particular, we apply random cropping, resizing and JPEG compression at different quality levels.
-More details on the dataset can be found in the supplemental material, here we only show a few examples in Fig.
+For example, when testing synthetic images generated by ProGAN, we use as real counterparts images from the LSUN dataset on which the ProGAN model was trained.
+The StyleGAN-T [64], eDiff-I [3], and GigaGAN [40] images were provided by the authors of the respective papers.
+Instead, we generated the images ourselves for Score-SDE [69], Stable Diff.
+[63], DiT [58], and Deepfloyd-IF [45] using the pre-trained models available online.
+Additional generated images were sourced from two public datasets [73, 14].
+For the text-driven commercial tools, we used a recently proposed dataset [4].
+Pristine and synthetic images with similar semantic content are built by extracting textual descriptions from a real dataset (RAISE [17]) and using them as prompts for generating synthetic images.
+A few examples of such images are shown in Fig.
 2.
+Overall, we have a dataset with $32,000$ real and fake images for all upcoming tests.
+To consider a more realistic scenario we simulate images shared on social networks and subject to post-processing operations: random cropping that can vary in a range from $\frac{5}{8}$ to the full size of the image, resizing to $200\times 200$ pixels, and JPEG compression with a random quality factor between $65$ and $100$.
 To measure performance, we consider three metrics.
-The area under the receiver-operating curve (AUC), and the average precision (AP), equal to the area under the precision-recall curve, are both threshold-independent.
+The area under the receiver operating curve (AUC) and average precision (AP) are both threshold-independent.
 The Accuracy instead, given by the number of correct predictions divided by the number of tested images, depends on the decision threshold.
-We use always a fixed threshold equal to 0.5, which fits a realistic scenario with no prior information on the data under test to carry out proper calibration.
+We always use a fixed threshold of 0.5, to simulate a realistic scenario in which no prior information on the data under test is available to carry out calibration.
 Generator
@@
 StyleGAN2 [43] u $256^{2}$-$1024^{2}$
-StyleGAN3 [43] u $256^{2}$-$1024^{2}$
+StyleGAN3 [44] u $1024^{2}$
 StyleGAN-T [64] t $512^{2}$
@@
 Adobe Firefly [31] t $2032^{2}$-$2048^{2}$
-Table 1: Image generators used in our experiments: GAN-based, DM-based, and commercial tools with models not always disclosed.
-For unconditional (u) and conditional (c) models, to prevent biases, the real images used in test come from the same datasets used to train the generator.
-Image resolution varies from 256$\times$256 to 2048$\times$2048.
+Table 1: Image generators used in our experiments: GAN-based, DM-based, and commercial tools.
+For unconditional (u) and conditional (c) models, the real images used in test come from the same datasets used to train the generator, while for text-to-image (t) models synthetic images have been generated using the prompt extracted from the real counterpart.
 4 CLIP for synthetic image detection
-We propose a simple procedure to tell apart real images from synthetic images based on features extracted from the image encoder of CLIP ViT L/14.
+We propose a simple procedure to distinguish real images from synthetic images based on features extracted from the image encoder of CLIP ViT L/14.
 The design of the detector consists of the following four steps:
@@
     use the captions to feed a text-driven image generator, $G(\cdot)$, so as to obtain $N$ synthetic images $\{F_{1},\ldots,F_{N}\}$ with $F_{i}=G(t_{i})$.
-We have now $N$ real / fake pairs that share the same textual description;
+Now we have $N$ real / fake pairs that share the same textual description;
   3. 3.
@@
   4. 4.
-    use the $N$+$N$ feature vectors to design a linear SVM classifier.
-If the real images have low-quality associated captions, or no caption at all, these can be easily generated by a dedicated tool such as BLIP [45].
+    use these $N$+$N$ vectors to design a linear SVM classifier.
+If the real images have low-quality associated captions, or no caption at all, these can be generated by a dedicated tool such as BLIP [46].
 Real and fake images are coupled to avoid possible semantic biases, which proves beneficial when the number of images is very small, e.g., $N=10$.
 We chose to extract features from the next-to-last layer rather than the last layer based on the results of preliminary experiments and, in the same way, we selected SVM among a set of candidate simple classifiers.
-All these preliminary experiments and their results are described in the supplemental material.
+All these preliminary experiments and their results are described in the supplementary material.
 In the following, we analyze the performance of the proposed detector as a function of the main design choices and experimental conditions.
 4.1 Influence of the reference set size
-We begin with investigating the role of the most important parameter of the proposed detector, the number of real / synthetic samples $N$ in the reference set, that in this experiment includes COCO real images and Latent Diffusion generated ones.
+We start by investigating the role of the most important parameter of the proposed detector, the number of real / synthetic samples $N$ in the reference set, which in this experiment includes COCO real images and Latent Diffusion synthetic images.
 Fig.
-3 shows results for $N$ varying from 10 to 100k in log scale.
-The results are presented in terms of AUC, AP, and Accuracy.
-When $N$ is less than 10k, we perform several runs with different sets of images and report the average metrics, e.g.
-for $N$ equals to 1k, we show the average on 10 runs.
-Let us focus first on the top row, showing results on images that have been neither re-compressed nor resized.
-In this simpler scenario, AUC and AP are extremely high, always above 85% for all known families of synthetic images (GAN, Diffusion) and somewhat worse for images generated by commercial tools.
-Especially remarkable is the fact that very good results are obtained already with as few as 10+10 reference images, showing this lightweight solution to be fully viable.
-The performance improves with $N$, growing by up to 5-10 percent points at the $N=10k$ plateau.
-We observe, in passing, that AUC and AP curves are very similar, as always happens when they approach 100%, hence we drop the latter for brevity in further experiments.
+3 shows results for $N$ ranging from 10 to 100k in log scale.
+Results are in terms of AUC, AP and Accuracy.
+When $N$ is less than 10k, we perform several runs with different sets of images and report the average metrics, e.g., for $N$ equals 1k, we average on 10 runs.
+Let’s focus first on the top row, which shows results on images that were neither recompressed nor resized.
+In this simpler scenario, both AUC and AP values are extremely high, consistently above 85% for all known families of synthetic images (GAN, Diffusion) and significantly worse only for images generated by commercial tools.
+It’s particularly remarkable that very good results are achieved with as few as 10+10 reference images, proving that this lightweight solution is fully viable.
+The performance improves with $N$, growing up to 5-10% at the 10k plateau.
+We note, in passing, that AUC and AP curves are very similar, so we drop the latter for brevity in further experiments.
 Figure 3: Performance of the CLIP-based detector as a function of the number of real and synthetic images in the reference set.
 We show AUC, AP and Accuracy on the original dataset (Top) and on post-processed images that simulate a realistic scenario (Bottom).
-In the rightmost figure, we see that unlike AUC and AP, accuracy does not always improve with increasing $N$.
-This is because AUC and AP are integral threshold-independent metrics.
-They tell us how well a method could do with a perfectly calibrated threshold.
-On the contrary, accuracy depends critically on the selected threshold which, lacking any prior knowledge, is set to 0.5.
-For Latent and GAN, 0.5 is a good threshold, and accuracy improves as the classifier does.
-Instead, for Diffusion and Commercial, the optimal threshold grows farther from 0.5 as $N$ increases and the accuracy decreases significantly.
-So there is a trade-off between optimality and robustness in our truly OOD scenario, where no calibration data exist (it is very unlikely that we know exactly which is the generator of the data under test).
-On the other hand, if we had data from the target class, even just 10 images, we could use them to design and ad hoc classifier.
+In the rightmost figure, we see that, unlike AUC and AP, accuracy does not always improve as $N$ increases.
+This is because AUC and AP are threshold-independent integral metrics.
+They tell us how well a method with a perfectly calibrated threshold might work.
+Instead, the precision depends critically on the selected threshold which, in the absence of any prior knowledge, is set at 0.5.
+For Latent, 0.5 is a good threshold and accuracy improves as $N$ grows.
+Instead, in general, the optimal threshold moves further away from 0.5 as $N$ increases and the precision decreases significantly.
+So, there is a trade-off between optimality and robustness in our truly OOD scenario, where no calibration data exists.
+On the other hand, if we had data from the target class, even just 10 images, we could use them to design an ad hoc classifier (see supplementary material).
 The bottom row of the figure shows results for images that have been compressed and/or resized.
-As expected, performance degrades somewhat in this scenario, but continues to be very good, with AUC between 75% and 90% and accuracy between 65% and 80% at $N$=10k.
-These are excellent results, much beyond the current state of the art, as we will show later.
-In this case, using only 10+10 images seems not to be advisable but the performance is near-optimal already at $N$=100.
-Datasets
-augm.
-GAN
-Diffusion
-Commerc.
-AVG
-family
-family
-tools
-COCO $+$ Latent
-93.9
-93.3
-81.7
-89.6
-COCO $+$ Latent
-84.8
-91.0
-87.6
-87.8
-COCO $+$ ProGAN
-95.0
-90.0
-65.6
-83.5
-COCO $+$ ProGAN
-92.3
-90.1
-82.3
-88.2
-LSUN $+$ Latent
-87.0
-82.7
-65.6
-78.5
-LSUN $+$ Latent
-86.8
-79.9
-80.3
-82.4
-LSUN $+$ ProGAN
-93.8
-79.2
-44.7
-72.6
-LSUN $+$ ProGAN
-95.7
-82.7
-67.5
-82.0
-Table 2: Performance of the CLIP-based detector for various real/fake combinations, with and without augmentation.
+As expected, performance degrades slightly in this scenario, but continues to be very good, with AUC between 75% and 90% and accuracy between 65% and 80% at $N$=10k.
+These are excellent results, well beyond the current state of the art, as we will show later.
+In this case, using only 10+10 images does not seem advisable but the performance is almost optimal already at $N$=100.
+Datasets augm. GAN Diffusion Commerc. AVG
+family family tools
+COCO $+$ Latent 92.4 92.6 80.5 88.5
+COCO $+$ Latent 89.3 91.8 87.0 89.4
+COCO $+$ ProGAN 93.6 90.0 65.6 83.1
+COCO $+$ ProGAN 91.3 90.1 82.3 87.9
+LSUN $+$ Latent 88.7 82.7 65.6 79.0
+LSUN $+$ Latent 87.9 79.9 80.3 82.7
+LSUN $+$ ProGAN 94.1 79.2 44.7 72.7
+LSUN $+$ ProGAN 95.0 82.7 67.5 81.7
+Table 2: AUC performance of the CLIP-based detector for various real/fake data, with and w/o augmentation (resizing/compression).
 4.2 Influence of the reference set content
-We discovered that not only the size but also the quality of the reference set images impacts heavily on performance.
-This probably does not need proof if we think about a very small reference set, like 10+10 images, where low quality or low diversity would spell catastrophe.
-However, something similar happens also when a larger reference set is used.
+We found that not only the quantity but also the quality of images in the reference set significantly impacts performance.
+This is intuitive with a very small reference set, like 10+10 images, where low quality or limited diversity could have catastrophic consequences.
+However, similar effects are observed even with a larger reference set.
 Tab.
 2 shows the results in terms of AUC obtained using various combination of real and synthetic images in the reference set, that is, (Real, Synth) $\in\{{\rm COCO,LSUN}\}\times\{{\rm Latent,ProGAN}\}$, with and without augmentation.
-In all cases, the reference set has size 10000+10000.
-Results are clearly affected by the specific combination.
-Indeed, there is a marked impairment when the LSUN dataset is used instead of COCO to draw the real samples and also, to a lesser extent, when ProGAN is used instead of Latent for the synthetic images.
-We conjecture that both LSUN and ProGAN are not diverse enough to describe adequately the decision domain.
-Moreover, LSUN has several polarizations not included in COCO, the images of LSUN represent only few well-defined categories, have all the same size, and are compressed most of them with the same quality factor.
+In all cases, the reference set has size 10,000+10,000.
+Results are clearly influenced by the specific combination.
+Indeed, considerable degradation occurs when the LSUN dataset is used instead of COCO to draw the real samples and, to a lesser extent, when using ProGAN instead of Latent for the synthetic images.
+We conjecture that both LSUN and ProGAN lack the diversity necessary to adequately describe the decision domain.
+Furthermore, unlike COCO, the LSUN dataset exhibits several biases: the images represent only a few well-defined categories, have all the same size, and most of them are compressed with the same quality factor.
 In Tab.
-2, we see that data augmentation only marginally reduces the effects of the used dataset.
+2, we see that data augmentation only marginally reduces the effects of the dataset used.
 Figure 4: Performance of the CLIP-based detector as a function of the pre-training.
@@
 Each curve refers to a different value of $N$ and we consider only the case of post processed images for brevity.
-For both metrics and all values of $N$ there is a steady increase in performance as larger and larger datasets are used for pre-training, from LAION-400M (0.4B images) [64], to LAION (2B) [65], to CommonPool (12.8B) [32].
-Overall, from smallest to largest dataset there is a ten point improvement!
-This comes to confirm, if ever necessary, the importance of pre-training a large VLM on the largest possible collection of varied images to acquire greater descriptive power for all kinds of applications.
-4.4 Beyond low-level forensic traces
-As already highlighted in the Introduction, forensic classifiers typically rely on subtle low-level artifacts.
-These can be seen more clearly in the frequency domain, where they appear as strong peaks.
-In this subsection we report on some initial analysis aimed at understanding whether the CLIP-based detector also relies on this type of traces or it exploits different information.
-To this end, we consider the high resolution 1024$\times$1024-pixel images generated by Stable Diffusion XL and resize them to 256$\times$256.
+For both metrics and all values of $N$ there is a steady increase in performance as increasingly larger datasets are used for pre-training, from LAION-400M (0.4B images) [65], to LAION (2B) [66], to CommonPool (12.8B) [33].
+Overall, there is a ten-point improvement from the smallest to the largest dataset.
+This confirms the importance of pre-training a large VLM on the largest possible collection of different images.
+Method Real/Synth. Training Size (k) Aug. Test Strategy
+[73] Wang et al. LSUN / ProGAN 360 / 360 global pooling
+[9] PatchFor. CelebA,FF / various 84 / 272 resizing
+[34] Grag. et al. LSUN / ProGAN 360 / 360 global pooling
+[49] Mand. et al. various / various 232 / 386 patch aggregation
+[48] Liu et al. LSUN / ProGAN 360 / 360 global pooling
+[14] Corvi et al. COCO,LSUN / Latent 180 / 180 global pooling
+[70] LGrad LSUN / ProGAN 72 / 72 resizing
+[57] Ojha et al. LSUN / ProGAN 360 / 360 cropping
+[74] DIRE-1 LSUN-Bed / ADM 40 / 40 resizing
+[74] DIRE-2 LSUN-Bed / St.GAN 40 / 40 resizing
+[71] NPR LSUN / ProGAN 72 / 72 resizing
+Ours 1k COCO / Latent 1 / 1 resizing
+Ours 1k+ COCO / Latent 1 / 1 resizing
+Ours 10k COCO / Latent 10 / 10 resizing
+Ours 10k+ COCO / Latent 10 / 10 resizing
+Table 3: List of methods. For each method, we report the datasets of real and synthetic images used for training, their sizes, whether or not augmentation is used, and the testing strategy.
+GAN family Diffusion family Commercial tools AVG
+Pro Style Style Style Giga Score ADM GLIDE Latent Stable DeepFl. Ediff-I DiT SDXL DALL· DALL· Midj. Adobe
+Method GAN GAN2 GAN3 GANT GAN SDE Diff. Diff. IF E2 E3 Firef.
+Wang et al. 100. 96.5 98.5 98.9 66.6 32.9 64.3 48.5 59.2 41.5 78.0 64.9 58.6 54.3 64.8 10.9 40.2 84.8 64.6
+PatchFor. 92.3 84.5 91.8 91.2 64.7 83.3 74.8 96.2 78.1 62.4 62.7 78.7 83.1 68.4 41.9 52.7 57.8 49.4 73.0
+Grag. et al. 100. 99.8 97.5 98.8 82.8 92.1 74.7 62.8 91.9 52.5 69.9 69.6 65.3 58.0 58.3  2.4 43.1 63.5 71.3
+Mand. et al. 96.2 93.8 100. 92.6 61.8 99.8 56.5 40.5 70.0 36.8 47.2 65.0 59.1 27.0 14.5 14.7 24.3 36.7 57.6
+Liu et al. 100. 99.8 98.4 98.5 98.2 95.4 82.5 76.5 97.6 77.4 72.2 98.7 88.0 31.1 70.4  0.2 40.7 11.8 74.3
+Corvi et al. 79.4 73.7 50.0 97.1 63.4 65.0 80.7 91.9 100. 100. 99.9 85.7 100. 100. 69.4 60.8 100. 98.0 84.2
+LGrad 100. 91.2 83.8 81.8 82.2 80.6 76.9 66.1 81.1 61.5 68.8 74.1 56.2 57.2 58.6 37.9 56.3 40.6 69.7
+Ojha et al. 100. 93.9 92.3 98.2 96.0 58.4 86.7 80.8 85.7 89.5 92.9 80.6 77.8 85.1 95.2 36.4 66.2 97.5 84.1
+DIRE-1 50.6 56.9 47.8 99.9 74.1 44.3 75.7 71.4 68.7 39.4 98.9 99.1 99.6 47.1 44.7 47.6 51.0 57.4 65.2
+DIRE-2 54.2 52.5 43.0 99.6 76.0 41.0 70.1 70.1 69.3 46.9 97.0 98.2 98.3 42.8 41.0 49.6 47.8 43.0 63.3
+NPR 100. 85.6 77.0 96.4 88.7 91.1 86.3 79.3 90.2 64.5 91.6 80.1 78.4 76.7 39.5 48.7 77.0 32.1 76.8
+Ours 1k 98.9 90.5 85.5 100. 81.3 89.1 81.1 99.9 94.1 87.6 96.5 98.5 94.1 87.8 89.0 70.0 73.0 74.4 88.4
+Ours 1k+ 91.4 80.9 84.0 99.8 74.7 84.3 75.2 99.6 81.6 89.8 98.0 99.1 92.5 88.9 83.6 93.6 78.7 85.1 87.8
+Ours 10k 99.8 91.8 86.8 100. 83.6 89.0 81.4 99.9 94.2 90.7 97.0 98.7 95.0 87.4 89.2 77.6 75.3 80.1 89.8
+Ours 10k+ 93.4 87.1 87.6 99.9 78.5 89.2 79.9 99.7 84.7 91.3 97.9 99.4 94.0 90.1 86.3 92.9 81.7 87.2 90.0
+Table 4: Comparison with SoTA methods in terms of AUC.
+For our approach we show four variants: 1k and 10k indicate the number of real and fake training images, + indicates augmentation (compression/resizing).
+Results on the dataset used for training are in light gray, while bold underlines the best performance for each dataset with a margin of 1%.
+The last column shows the average over all datasets.
+GAN family Diffusion family Commercial tools AVG
+Pro Style Style Style Giga Score ADM GLIDE Latent Stable DeepFl. Ediff-I DiT SDXL DALL· DALL· Midj. Adobe
+Method GAN GAN2 GAN3 GANT GAN SDE Diff. Diff. IF E2 E3 Firef.
+Wang et al. 100. 86.6 88.4 61.7 59.2 68.0 65.0 60.6 67.1 55.2 50.3 48.3 55.1 64.5 46.2 27.7 46.7 55.9 61.5
+PatchFor. 57.9 51.8 57.0 50.6 53.8 69.0 66.2 83.3 58.8 48.3 61.3 65.0 68.1 63.3 64.3 63.3 59.0 65.1 61.4
+Grag. et al. 100. 95.4 90.9 94.4 64.4 77.1 77.1 79.8 84.0 53.5 50.6 55.6 66.7 66.6 55.2 25.1 48.5 60.2 69.2
+Mand. et al. 81.1 79.3 87.2 49.1 49.3 64.0 54.8 42.6 52.9 39.4 55.7 54.5 49.8 42.2 47.9 42.3 35.2 53.4 54.5
+Liu et al. 64.3 55.1 50.1 57.3 45.4 62.6 51.1 58.6 50.7 58.6 50.9 64.2 53.9 56.0 44.4 61.8 52.6 53.1 55.0
+Corvi et al. 77.5 74.7 69.4 82.1 66.6 70.4 79.0 93.5 99.3 69.9 60.7 72.1 89.2 61.8 65.9 32.4 51.9 58.1 70.8
+LGrad 56.3 58.3 49.8 52.3 43.5 45.9 49.2 42.3 50.4 54.8 40.7 46.4 49.4 53.2 41.8 53.5 50.4 51.8 49.4
+Ojha et al. 99.8 75.5 75.4 91.1 88.5 79.3 83.7 83.3 81.8 75.0 59.9 68.7 70.1 61.8 63.2 41.7 40.6 52.9 71.8
+DIRE-1 48.4 42.5 39.1 53.5 54.3 44.1 48.0 44.7 46.1 47.0 66.2 62.8 53.2 47.1 44.6 47.6 51.0 57.4 49.9
+DIRE-2 49.3 41.6 38.6 53.8 55.0 44.3 45.1 40.2 45.9 56.4 70.7 72.2 53.0 42.8 40.9 49.7 47.8 43.0 49.5
+NPR 54.5 48.5 41.9 54.0 44.8 44.7 46.9 47.2 47.7 55.4 49.6 54.6 50.9 52.8 50.0 67.5 50.8 55.5 51.0
+Ours 1k 85.0 64.0 66.6 90.2 75.2 74.7 78.1 97.2 77.1 77.6 80.1 86.6 77.5 76.5 77.9 77.4 63.1 70.5 77.5
+Ours 1k+ 78.7 62.5 68.4 97.5 67.9 84.0 74.3 99.6 78.2 83.7 94.5 97.1 88.9 89.6 81.2 90.9 77.6 83.7 83.2
+Ours 10k 85.7 65.5 68.1 90.5 74.7 75.8 78.4 97.7 77.8 78.1 81.2 87.1 77.2 76.4 78.2 76.4 65.0 72.2 78.1
+Ours 10k+ 82.8 67.4 70.7 98.4 71.9 85.4 77.3 99.7 80.2 85.8 95.9 98.2 91.1 89.9 83.8 90.1 79.4 85.5 85.2
+Table 5: Comparison with SoTA methods (AUC) on images post-processed by random cropping, resizing and compression.
+5 Comparison with the state-of-the-art
+In this Section we perform an extensive comparison with SoTA methods in different scenarios.
+We consider four versions of our approach, with 1,000+1,000 (1k) or 10,000+10,000 (10k) real+fake images in the reference set, and with (+) or without compressed/resized images for augmentation.
+To ensure a fair comparison, we only include SoTA methods with code and/or pre-trained models publicly available online.
+They are listed in Tab.
+3 and described in the supplementary material.
+Generalization analysis.
+In Tab.
+4 we show the results in terms of AUC on 18 generative models.
+For each method, the results obtained on the same dataset used for training are in light gray, since the ID scenario is of little interest for our analysis.
+The items in bold, instead, highlight the best OOD performance for each dataset, considering a margin of 1%.
+We can observe that methods trained on one GAN dataset generally perform well on other datasets of the same family but fare much worse on DM datasets.
+Then, the roles change for methods trained on DM datasets.
+This is unsurprising, as generators of the same family share architectural details that leave similar traces on the generated images.
+No SoTA method performs uniformly well on all datasets.
+In contrast, the proposed lightweight CLIP-based detector consistently delivers strong performance.
+The version with 10k+10k reference images, without augmentation, outperforms the best competitor by +6.8% in terms of average AUC.
+For GAN-based generators, the proposed CLIP-based detectors keep providing the best performance, generally much better than SoTA methods, except for the Liu method which works almost at the same level.
+However, the performance of this method, as well as many other methods, decreases catastrophically when considering synthetic images from commercial tools.
+This is the most realistic and interesting scenario, with images of unknown origin and no prior information on the possible generation process.
+In this situation, most methods provide unreliable decisions.
+Corvi’s method works surprisingly well on some datasets, perhaps generated by diffusion models similar to Latent.
+However, only the proposed CLIP-based approach provides good stable performance across all cases.
+On commercial tools, the versions with resized and recompressed images in the reference set prove especially strong.
+Robustness to perturbations.
+The observed trend with respect to unknown models is further accentuated when the images undergo post-processing, as shown in Tab.
+5.
+These impairments attenuate forensic traces, to the point that most SoTA methods become essentially useless, performing no better than random chance, particularly on unknown commercial models.
+CLIP-based detectors, instead, keep providing a good performance.
+6 Beyond low-level forensic traces
+As previously mentioned, conventional detectors often experience significant performance degradation when images are resized or compressed, as the subtle forensic traces they depend on are significantly reduced.
+In contrast, the proposed CLIP-based detector keeps working well under these conditions, suggesting that it relies on higher-level semantic features.
+To further investigate this hypothesis, we now conduct some targeted experiments.
+Removing/inserting low-level traces in images.
+We take synthetic images generated by Stable Diffusion 1.4, Stable Diffusion 2.0 [63], Stable Diffusion XL [59]) and downsample them by 1/4.
 Fig.
-5 shows, on the right, the high and low-resolution images (top) together with their spectra (bottom).
-We observe that after resizing, the spectral artifacts have completely disappeared.
-In fact, these traces are caused by the aliasing effect of the upsampling filters, which can be removed by proper decimation.
-We run our detector on 1000 original and resized synthetic images, obtaining in both cases the same True Positive Rate (TPR) of $43\%$ (note that the True Negative Rate (TNR) for these images is $97\%$).
-We also carry out a complementary experiment where we take real images and pass them through the generative architecture of Stable Diffusion which acts as an autoencoder.
-Therefore, the output (Fig.
-5 right) is visually identical to the input but its Fourier spectrum now shows strong peaks.
-Even so, the CLIP-based detector is not affected by the presence of such traces and the TNR, $97\%$, is almost the same as with unprocessed images.
-We find this last experiment to be particularly relevant.
-In fact, it is very important to take into account malicious operations that can be carried out with the aim of reducing the credibility of current forensic detectors.
-Of course, several forms of attack can be applied, but this one seems to be especially simple and disruptive, such to easily fool detectors that rely exclusively on low-level traces.
-Based on this experiment, it seems that the proposed CLIP-based detector looks for different types of traces.
-Therefore, besides its stand-alone performance, it could be used to complement and enhance the performance of other detectors.
-A larger set of experiments can be found in the supplementary.
+5 shows, on the left, the high- and low-resolution images (top) along with their spectra (bottom) for Stable XL.
+After reducing the resolution, the peaks in the Fourier domain completely disappeared, because they were caused by the aliasing effect of the oversampling filters, which can be removed by appropriate decimation.
+In a complementary experiment, we take real images from the RAISE dataset and pass them through the autoencoders used by Stable Diffusion XL.
+This process does not change the visual appearance of the images (Fig.
+5 top-right) but inserts low-level forensic traces in their spectra (bottom) that resemble those of the generated images.
+In Table 6 we quantify the effects of such attacks on 1,000 real and 1,000 synthetic images, reporting the AUC for our method and the best competitor before the attack (first group of columns) after removing traces from synthetic images (second group) and after adding them to real images (third group).
+Before the attack, all detectors work quite well, with the method proposed by Corvi et al., based on low-level features, reaching perfect discrimination.
+After the attacks, however, the performance of Corvi et al.
+degrades dramatically, while the proposed method suffers only a very limited loss.
+In summary, the CLIP-based detector appears to withstand various types of image degradations easily, both innocent and malicious.
+The latter case is particularly relevant: attacks can be carried out with the very aim of making everything appear as synthetic, misleading current forensic detectors based on low-level traces, and undermining public trust in forensic analyses.
+Fusion.
+Based on the above results, the proposed CLIP-based detector appears not to rely on the same low-level traces exploited by most of the current detectors.
+This is further supported by the scatter plot of Fig.6 where each point gives the scores of the CLIP-based and Corvi detectors for a given image.
+The two sets of scores are almost orthogonal, as if they depended on uncorrelated features, a property that paves the way for appropriate fusion strategies.
+We implemented a simple decision rule where the image is declared real only if both detectors agree on this choice.
+Tab.7 reports AUC and Accuracy results for Corvi, the proposed method (10k and 10k+ versions), and their fusion.
+Results are given as averages over the three families of generators: GAN, Diffusion and Commercial Tools.
+The fusion ensures a further boost in performance over the proposed CLIP-based detector, both in terms of AUC (+3.6%) and Accuracy (+7.4%).
+A smaller improvement is observed when images are resized/compressed, arguably because low-level traces are more compromised in this case.
 Refer to caption Refer to caption Refer to caption Refer to caption
 Refer to caption Refer to caption Refer to caption Refer to caption
-Figure 5: Top (from left to right): a high resolution synthetic image generated by Stable Diffusion XL [58] and its 4$\times$ decimated version; a real image and the corresponding image obtained at the output of Stable Diffusion [62] used as an autoencoder.
+Figure 5: Top (from left to right): a synthetic image generated by Stable Diffusion XL [59] and its 4$\times$ decimated version; a real image and the corresponding image processed by the autoencoder of Stable Diffusion XL.
 Bottom: Fourier spectra of the noise residuals for images shown on the top.
 A suitable decimation removes Fourier peaks in synthetic images, while passing a real image through an autoencoder creates new peaks.
-GAN family Diffusion family AVG
-Pro
-Style
-Style
-Style
-Giga
-Score
-ADM GLIDE
-Latent
-Stable
-DeepFl.
-Ediff-I DiT SDXL
-Method
-real/synth. training
-size (k)
-GAN
-GAN2
-GAN3
-GANT
-GAN
-SDE
-Diff.
-Diff.
-IF
-Wang et al.
-LSUN / ProGAN
-360 / 360
-100.
-96.5
-89.0
-98.9
-66.6
-32.9
-64.3
-48.5
-59.2
-41.5
-78.0
-64.9
-58.6
-54.3
-68.1
-PatchFor.
-CelebA,FF / various
-84 / 272
-92.3
-84.5
-87.9
-91.2
-64.7
-83.3
-74.8
-96.2
-78.1
-62.4
-62.7
-78.7
-83.1
-68.4
-79.2
-Grag. et al.
-LSUN / ProGAN
-360 / 360
-100.
-99.8
-98.3
-98.8
-82.8
-92.1
-74.7
-62.8
-91.9
-52.5
-69.9
-69.6
-65.3
-58.0
-79.8
-Liu et al.
-LSUN / ProGAN
-360 / 360
-100. 99.8
-97.8
-98.5
-98.2 95.4
-82.5
-76.5
-97.6
-77.4
-72.2
-98.7
-88.0
-31.1
-86.7
-Corvi et al.
-COCO,LSUN / Latent
-180 / 180
-79.4
-73.7
-38.4
-97.1
-39.9
-65.0
-80.7
-91.9
-100. 99.7 99.9
-85.7
-100. 100.
-82.2
-LGrad
-LSUN / ProGAN
-72 / 72
-100.
-91.2
-87.3
-81.8
-82.2
-80.6
-76.9
-66.1
-81.1
-61.5
-68.8
-74.1
-56.2
-57.2
-76.1
-Ojha et al.
-LSUN / ProGAN
-360 / 360
-100.
-93.9
-90.9
-98.2
-96.0
-58.4
-86.7
-80.8
-85.7
-89.5
-92.9
-80.6
-77.8
-85.1
-86.9
-DIRE-1
-LSUN-Bed / ADM
-40 / 40
-50.6
-56.9
-49.3
-99.9
-74.1
-44.3
-75.7
-71.4
-68.7
-39.4
-98.9
-99.1
-99.6
-47.1
-69.7
-DIRE-2
-LSUN-Bed / St.GAN
-40 / 40
-54.2
-52.5
-44.3
-99.6
-76.0
-41.0
-70.1
-70.1
-69.3
-46.9
-97.0
-98.2
-98.3
-42.8
-68.6
-Ours 1k
-COCO / Latent
-1 / 1
-98.7
-89.8
-91.9
-100.
-81.2
-89.4
-80.3
-99.9 94.0
-87.4
-96.8
-98.7
-93.9
-89.2
-92.2
-Ours 1k+
-COCO / Latent
-1 / 1
-89.2
-79.7
-83.2
-99.7
-74.2
-84.0
-74.9
-99.6
-82.0
-89.5
-97.8
-99.2
-91.8
-90.3
-88.2
-Ours 10k
-COCO / Latent
-10 / 10
-99.8
-92.5
-92.7
-100.
-84.4
-88.0
-83.2
-99.9
-96.2
-91.8
-97.6
-99.2
-95.2
-88.8
-93.5
-Ours 10k+
-COCO / Latent
-10 / 10
-87.1
-81.8
-81.9
-99.8
-73.5
-85.7
-79.1
-99.3
-82.5
-90.6
-98.4
-99.7
-93.2
-90.8
-88.8
-Table 3: Comparison with SOTA methods in terms of AUC.
-For our approach we show four variants: trained on 1k real and 1k fake images; 10k real and 10k fake images; trained on 1k and 1k fake images but including compressed/resized images (1k+) and trained on 10k and 10k fake images but including compressed/resized images (10k+).
-Results obtained on the very same dataset used for training are in light gray, while entries in bold underline the best OOD performance for each dataset.
-We also report the mean AUC by averaging the AUC scores over all datasets.
-GAN family Diffusion family AVG
-Pro
-Style
-Style
-Style
-Giga
-Score
-ADM GLIDE
-Latent
-Stable
-DeepFl.
-Ediff-I DiT SDXL
-Method
-real/synth. training
-size (k)
-GAN
-GAN2
-GAN3
-GANT
-GAN
-SDE
-Diff.
-Diff.
-IF
-Wang et al.
-LSUN / ProGAN
-360 / 360
-100.
-69.2
-56.2
-71.5
-52.8
-50.4
-50.3
-50.7
-51.3
-50.1
-50.0
-50.4
-50.1
-49.8
-57.3
-PatchFor.
-CelebA,FF / various
-84 / 272
-64.8
-50.9
-50.1
-51.1
-51.3
-50.0
-50.3
-50.5
-51.3
-50.2
-50.0
-50.0
-51.2
-50.0
-51.5
-Grag. et al.
-LSUN / ProGAN
-360 / 360
-100. 96.6 92.6
-91.9
-69.8
-83.8
-54.5
-56.8
-67.5
-48.9
-48.9
-52.1
-49.2
-51.4
-68.9
-Liu et al.
-LSUN / ProGAN
-360 / 360
-100.
-96.2
-78.6
-89.4
-93.3 89.2
-67.2
-73.2
-84.3
-65.3
-50.5
-92.0
-64.3
-39.6
-77.4
-Corvi et al.
-COCO,LSUN / Latent
-180 / 180
-50.0
-50.7
-50.0
-52.0
-57.7
-50.0
-54.5
-57.3
-99.8 99.7
-64.3
-50.9
-99.8 99.7
-66.9
-LGrad
-LSUN / ProGAN
-72 / 72
-100.
-77.2
-62.3
-68.5
-74.7
-60.3
-61.2
-55.4
-69.3
-61.5
-60.9
-62.6
-53.2
-58.7
-66.1
-Ojha et al.
-LSUN / ProGAN
-360 / 360
-99.9
-77.3
-82.4
-87.1
-81.2
-57.1
-60.9
-57.1
-68.1
-72.6
-65.7
-66.0
-55.4
-68.0
-71.3
-DIRE-1
-LSUN-Bed / ADM
-40 / 40
-51.4
-50.8
-50.0
-96.5
-73.1
-50.0
-73.2
-73.2
-73.3
-49.7
-95.3 95.6
-94.9
-49.9
-69.8
-DIRE-2
-LSUN-Bed / St.GAN
-40 / 40
-51.4
-50.7
-50.0
-82.0
-66.6
-50.0
-65.9
-65.9
-66.4
-49.6
-80.7
-81.5
-82.2
-49.6
-63.7
-Ours 1k
-COCO / Latent
-1 / 1
-69.7
-72.9
-82.0
-98.9
-69.0
-76.6
-71.2
-97.8 86.9
-70.9
-63.2
-80.3
-81.0
-60.2
-77.2
-Ours 1k+
-COCO / Latent
-1 / 1
-68.4
-70.3
-73.5
-96.2
-62.9
-72.8
-65.8
-94.7
-74.5
-77.1
-81.6
-91.9
-78.0
-75.2
-77.4
-Ours 10k
-COCO / Latent
-10 / 10
-90.1
-83.0
-84.0
-98.7
-66.5
-75.6
-72.3
-93.7
-89.9
-72.9
-59.5
-78.8
-76.8
-54.4
-78.3
-Ours 10k+
-COCO / Latent
-10 / 10
-73.8
-72.3
-71.4
-95.0
-60.8
-75.6
-68.2
-90.5
-74.4
-77.3
-81.7
-92.4
-77.4
-70.0
-77.2
-Table 4: Comparison with SOTA methods in terms of Accuracy, fixed 0.5 threshold.
-Results obtained on the very same dataset used for training are in light gray, while entries in bold underline the best OOD performance for each dataset.
-We also report the mean Accuracy by averaging the Accuracy scores over all datasets.
-5 Comparison with the state-of-the-art
-In this Section we perform an extensive comparison with SoTA methods in different scenarios.
-For our approach we show four variants: trained on 1k real and 1k fake images; 10k real and 10k fake images; trained on 1k and 1k fake images but including compressed/resized images (1k+) and trained on 10k and 10k fake images but including compressed/resized images (10k+).
-In our analysis we only include SoTA techniques with code and/or pre-trained models publicly available on-line in order to ensure a fair comparison.
-We considered the following approaches:
+Refer to captionRefer to caption
+Figure 6: Scatter plots of scores provided by Corvi et al. ($y$-axis) and proposed method without and with augmentation ($x$-axis) on 2400+2400 images sampled from the 18 datasets of Tab.1.
+7 Discussion
+We have proposed a simple method based on CLIP features to distinguish real images from synthetic images. Through extensive experimental analysis, we found that:
+  • CLIP features support a much higher generalization ability than what had been discovered so far.
+By leveraging just a few examples, not even belonging to the generator under test, a simple CLIP-based detector achieves top performance on a large variety of generators and in the most challenging conditions.
+  • Maximizing the diversity of reference CLIP features has a positive impact on performance, even when a relatively large number of examples is considered.
+  • It is known that CLIP’s descriptive power comes from its huge pre-training set. Further increasing this set keeps boosting performance, up to $10\%$.
+  • Experimental evidence suggests that CLIP features, even when adapted to forensic applications, are largely independent of low-level forensic traces.
+This provides some immunity to malicious attacks targeting low-level artifacts and paves the way for further performance improvements through suitable fusion with traditional detectors.
+The present study provides a number of interesting insights but leaves much room for future work.
+We believe that new and better forensic methods can be proposed based on CLIP features.
+One major area of study is the development of few-shot methods that adapt the detector to the situation of interest on the fly.
+Future work should also consider interpretability, starting with understanding which forensic features the detector exploits to make its decisions.
+downsample by 1/4 add low-level traces
+Method St.Diff. St.Diff. SDXL St.Diff. St.Diff. SDXL St.Diff. St.Diff. SDXL
+1.4 2.0 1.4 2.0 1.4 2.0
+Corvi et al. 100. 100. 100. 41.6 46.2 45.6 79.3 80.4 67.6
+Ours 1k 88.8 87.0 87.8 85.8 82.2 79.3 89.3 82.5 80.8
+Ours 1k+ 90.2 89.4 88.9 89.8 91.1 90.0 90.8 86.1 83.2
+Ours 10k 93.8 90.6 87.4 86.8 84.4 80.5 93.7 86.2 80.6
+Ours 10k+ 94.0 90.4 90.1 95.9 92.6 91.6 94.1 86.7 84.0
+Table 6: Results in terms of AUC without attacks (first group of columns) and after attacks on synthetic images (second group) and real images (third group).
+Corvi et al., based on low-level forensic traces, is severely affected by the attacks while the proposed method keeps working well in all conditions.
+Families of Generators
+GAN Diffusion Comm. Tools Average
+Method AUC/Acc AUC/Acc AUC/Acc AUC/Acc
+Corvi et al. 72.7 / 52.1 91.5 / 75.1 82.1 / 62.8 82.1 / 63.3
+Ours 10k 92.4 / 79.1 92.6 / 73.3 80.5 / 52.6 88.5 / 68.3
+Ours 10k+ 89.3 / 74.9 91.8 / 77.2 87.0 / 67.3 89.4 / 73.1
+Ours fusion 92.9 / 80.1 96.9 / 87.8 88.2 / 65.3 92.7 / 77.8
+Ours fusion+ 89.9 / 75.5 96.8 / 88.8 92.3 / 77.2 93.0 / 80.5
+Corvi et al. 74.0 / 55.1 77.3 / 62.1 52.1 / 50.1 67.8 / 55.8
+Ours 10k 76.9 / 63.6 81.1 / 63.7 73.0 / 52.4 77.0 / 59.9
+Ours 10k+ 78.2 / 69.0 89.3 / 78.7 84.7 / 66.4 84.1 / 71.4
+Ours fusion 78.5 / 66.8 85.1 / 71.4 72.7 / 52.6 78.7 / 63.6
+Ours fusion+ 79.6 / 70.8 91.7 / 80.7 84.5 / 66.6 85.3 / 72.7
+Table 7: AUC/Accuracy results for Corvi et al., proposed method (10k and 10k+ versions), and their fusion over the three families of generators: GAN, Diffusion, Commercial Tools.
+Top: original images; bottom: compressed/resized images.
+Acknowledgment.
+We gratefully acknowledge the support of this research by a TUM-IAS Hans Fischer Senior Fellowship, a TUM-IAS Rudolf Mößbauer Fellowship and a Google Gift.
+This material is also based on research sponsored by the Defense Advanced Research Projects Agency (DARPA) and the Air Force Research Laboratory (AFRL) under agreement number FA8750-20-2-1004.
+In addition, this work has received funding by the European Union under the Horizon Europe vera.ai project, Grant Agreement number 101070093.
+Finally, we want to thank Koki Nagano, Tero Karras, Yogesh Balaji and Ming-Yu Liu for sharing data for StyleGAN-T and eDiff-I experiments.
+Supplemental Material
+In this appendix, we provide a brief description of the methods used for comparison (Sec.
+A), report additional ablation studies (Sec.
+B), additional results (Sec.
+C), and robustness analysis (Sec.
+D).
+We also show some experiments carried out on social networks in a few-shot scenario (Sec.
+E).
+Finally, we enlarge our initial dataset with additional synthetic generators and carry out further experiments on generalization (Sec.
+F).
+Appendix A Reference methods
+In our experimental analysis we included the following methods:
   1. (a)
@@
   2. (b)
-    Grag et al. [33] proposes a simple modification to the ResNet50 architecture which allows to better preserve low-level forensic traces and is trained on the same dataset introduced in [71].
+    Gragnaniello et al.
+[34] proposes a simple modification to the ResNet50 architecture which allows to better preserve low-level forensic traces and is trained on the same dataset introduced in [73].
   3. (c)
+    Mandelli et al.
+[49] relies on the ensemble of five EfficientNet-B4 networks trained on different datasets.
+At test-time the scores of randomly selected patches are aggregated: if at least one patch is detected as synthetic, then the entire image is classified as synthetic.
+  4. (d)
     PatchForensics [9] develops a fully-convolutional classifier based on local patches with limited receptive fields over an XceptionNet backbone.
-  4. (d)
+  5. (e)
     Liu et al.
@@
 This spatial information is combined with frequency information to improve the classification.
-  5. (e)
+  6. (f)
     LGrad [70] works on the gradients extracted through a pre-trained CNN model in order to filter out the content of the image and transform a data-dependent problem into a transformation-model dependent problem.
-  6. (f)
+  7. (g)
     Corvi et al. [14] performs strong augmentation to gain robustness and increase generalization and is trained on a large dataset of latent diffusion models.
-  7. (g)
-    Ojha et al. [56] uses CLIP-based features and perfoms detection using the CLIP feature space. Both nearest neighbor and linear probing are tested, even if the latter gives better performance.
   8. (h)
+    Ojha et al. [57] uses a large dataset of real and synthetic images to train a simple classifier working on pre-trained CLIP features.
+  9. (i)
     DIRE [74] is based on the idea that synthetic images can be reconstructed better than real images by a pre-trained model.
-To this end it is proposed to train a ResNet-50 in two different ways, using ADM images (DIRE 1) and StyleGAN images (DIRE 2), respectively.
-Comparison over GAN- and DM-based generators.
-In Tab.
-3 and Tab.
-4 we show results in terms of AUC and Accuracy, respectively, over 14 GAN-based and DM-based generators proposed in the scientific literature and perfectly known.
-For each method, results obtained on the very same dataset used for training are in light gray, to signal an ID scenario, of little interest for our analysis.
-Entries in bold, instead, underline the best OOD performance for each dataset.
-Focusing on AUC data, Tab.
-3, we observe that methods trained on a GAN dataset work generally well on other dataset of the same family, and much worse on DM datasets.
-Then, the roles switch for methods trained on DM datasets.
-This is nothing new, as generators of the two families share architectural details that leave similar traces on the generated images.
-No SoTA method works uniformly well on all datasets.
-On the contrary, the proposed lightweight CLIP-based detector provides a uniformly good performance, and the version with 10k+10k reference images without augmentation outperforms the best competitor (notably, again a CLIP-based method) by +6.6%, obtaining an average AUC well above 93.5% a truly impressive accomplishment.
-In terms of Accuracy, Tab.
-4, performance lowers significantly, and in some cases dramatically.
-The proposed CLIP-based detectors keep providing the best performance, generally much better than SoTA methods, except for Liu’s method that performs almost at the same level.
-Comparison over commercial tools.
-However, the performance of Liu’s method, as well as that of several other methods, drops catastrophically when considering the Commercial Tools (DALL·E2, DALL·E3, Midjourney, FireFly) both in terms oa AUC and Accuracy, see Tab.
-5.
-Needless to say, this is the most realistic and interesting scenario, often with images of unknown origin and no prior information on the possible generation process.
-In this situation, most methods provide unreliable decisions.
-Corvi’s method works surprisingly well on some datasets, maybe generated by diffusion models similar to Latent.
-However, only the proposed CLIP-based approach provides a stable good performance in all cases.
-Note that, now, the versions with resized and recompressed images in the reference set perform significantly better than those without augmentation.
-Robustness to perturbations.
-The trend observed over unknown models is further accentuated when images are subject to some post-processing, Tab.
-6.
-These impairments tend to wash out forensic traces, to the point that most SoTA methods become basically useless with coin-tossing performance, especially on the unknown commercial models.
-Again, the CLIP-based detectors provide a more than decent performance, further supporting our conjecture on the high-level nature of the forensic traces they rely upon.
-Commercial tools
-DALL·E2
-DALL·E3
-Midjourney
-FireFly
-Average
-Method
-AUC/Acc
-AUC/Acc
-AUC/Acc
-AUC/Acc
-AUC/Acc
-Wang et al.
-64.8 / 50.2
-10.9 / 49.4
-40.2 / 49.5
-84.8 / 50.3
-50.2 / 49.8
-PatchFor.
-41.9 / 50.0
-52.7 / 50.0
-57.8 / 50.0
-49.4 / 50.0
-50.5 / 50.0
-Grag. et al.
-58.3 / 49.2
-2.4 / 36.9
-43.1 / 40.5
-63.5 / 47.5
-41.8 / 43.6
-Liu et al.
-70.4 / 66.4
-0.2 / 30.8
-40.7 / 43.3
-11.8 / 31.0
-30.8 / 42.9
-Corvi et al.
-69.4 / 50.1
-60.8 / 50.0
-100. / 100.
-98.0 / 51.1
-82.1 / 62.8
-LGrad
-58.6 / 55.5
-37.9 / 47.8
-56.3 / 57.9
-40.6 / 47.5
-48.3 / 52.1
-Ojha et al.
-95.2 / 83.5
-36.4 / 47.3
-66.2 / 52.4
-97.5 / 89.9
-73.8 / 68.3
-DIRE-1
-44.7 / 50.0
-47.6 / 50.0
-51.0 / 49.7
-57.4 / 49.9
-50.2 / 49.9
-DIRE-2
-41.0 / 49.9
-49.6 / 50.0
-47.8 / 49.8
-43.0 / 49.8
-45.3 / 49.9
-Ours 1k
-89.1 / 62.2
-77.6 / 53.4
-74.3 / 53.2
-77.1 / 52.1
-79.5 / 55.2
-Ours 1k+
-83.9 / 68.2
-94.8 / 83.1
-79.7 / 64.7
-86.0 / 68.4
-86.1 / 71.1
-Ours 10k
-88.8 / 57.8
-82.9 / 51.1
-75.4 / 51.2
-79.9 / 50.9
-81.7 / 52.8
-Ours 10k+
-84.5 / 63.9
-94.4 / 76.6
-83.8 / 62.7
-87.6 / 64.5
-87.6 / 67.0
-Table 5: AUC/Accuracy on Commercial Tools. We show AUC and Accuracy on very recent synthetic generators: DALL·E2, DALL·E3, Midjourney v5 and Firefly.
-Families of Generators
-GAN
-Diffusion
-Comm. Tools
-Average
-Method
-AUC/Acc
-AUC/Acc
-AUC/Acc
-AUC/Acc
-Wang et al.
-78.9 / 62.1
-59.3 / 50.4
-44.1 / 49.9
-60.8 / 54.1
-PatchFor.
-54.3 / 50.0
-64.8 / 50.3
-62.9 / 50.1
-60.7 / 50.1
-Grag. et al.
-89.2 / 68.5
-67.9 / 50.8
-47.3 / 50.0
-68.1 / 56.4
-Liu et al.
-54.7 / 51.6
-56.3 / 51.1
-53.0 / 50.7
-54.6 / 51.1
-Corvi et al.
-66.7 / 54.1
-76.1 / 62.1
-52.1 / 50.1
-65.0 / 55.4
-LGrad
-51.2 / 51.0
-48.0 / 50.5
-49.4 / 50.4
-49.5 / 50.7
-Ojha et al.
-86.9 / 74.9
-73.7 / 59.8
-49.6 / 50.6
-70.1 / 61.8
-DIRE-1
-48.9 / 50.3
-51.0 / 50.5
-50.1 / 49.9
-50.0 / 50.2
-DIRE-2
-49.1 / 50.0
-52.3 / 52.5
-45.3 / 49.9
-48.9 / 50.8
-Ours 1k
-76.4 / 68.7
-81.2 / 71.4
-74.3 / 61.9
-77.3 / 67.3
-Ours 1k+
-75.3 / 67.8
-87.8 / 77.6
-84.2 / 66.0
-82.4 / 70.4
-Ours 10k
-76.8 / 60.7
-82.7 / 59.6
-73.8 / 50.5
-77.8 / 56.9
-Ours 10k+
-78.2 / 67.9
-88.7 / 77.0
-83.0 / 65.6
-83.3 / 70.2
-Table 6: AUC/Accuracy in the presence of post-processing. All the images have been randomly cropped, resized and compressed to simulate a realistic scenario over the web.
-6 Discussion
-In this work, we have shown a series of good features of CLIP that can help detecting real from synthetic content. Through an extensive experimental analysis we found that:
-  • CLIP features support a much higher generalization ability than what had been discovered so far.
-By leveraging just a few examples, not even belonging to the generator under test, a simple CLIP-based detector achieves top performance on a large variety of generators and in the most challenging conditions.
-  • Maximizing the diversity of reference CLIP features impacts positively on performance, even when a relatively large number of examples is considered.
-  • It is well known that CLIP descriptive power descends from its huge pre-training set. Further increasing this set keeps boosting the performance, up to $10\%$.
-  • Experimental evidence seems to show that CLIP features, even when tuned to forensic applications, are mostly independent from low-level forensic traces.
-This opens the way to interesting developments.
-On one hand, one can gain immunity to possible malicious attacks working on low-level artifacts.
-On the other hand, there is room for further performance improvement by suitable fusion with classic detectors.
-Overall, the present study provides a number of interesting insights but leaves much room for future work.
-We believe that new and better forensic methods can be proposed on top of CLIP.
-One major area of study can be to develop few-shot methods that adapt the detector on the fly to the situation of interest.
-Future work should also consider interpretability, to begin with understanding which forensic features is the detector exploiting to make its decisions.
-Acknowledgment.
-We gratefully acknowledge the support of this research by a TUM-IAS Hans Fischer Senior Fellowship, a TUM-IAS Rudolf Mößbauer Fellowship and a Google Gift.
-This material is also based on research sponsored by the Defense Advanced Research Projects Agency (DARPA) and the Air Force Research Laboratory (AFRL) under agreement number FA8750-20-2-1004.
-The U.S.
-Government is authorized to reproduce and distribute reprints for Governmental purposes notwithstanding any copyright notation thereon.
-The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either expressed or implied, of DARPA or the U.S.
-Government.
-In addition, this work has received funding by the European Union under the Horizon Europe vera.ai project, Grant Agreement number 101070093.
-It is also supported by the PREMIER project, funded by the Italian Ministry of Education, University, and Research within the PRIN 2017 program.
-Finally, we want to thank Koki Nagano for useful discussions, and Tero Karras, Yogesh Balaji, Ming-Yu Liu for sharing data for StyleGAN-T and eDiff-I experiments.
-Supplemental Material
-In this appendix, we report additional ablation studies (Sec.
-A) and robustness analysis (Sec.
-B).
-We also enlarge our initial dataset with additional synthetic generators and carry out further experiments on generalization (Sec.
-C).
-Furthermore, we conduct more experiments on the ability of our approach and SoTA methods to deal with possible attacks from malicious users (Sec.
-D).
-Finally, we show some experiments carried out on social networks in a few-shot scenario (Sec.
-E).
-Figure 6: Average AUC for our approach considering different classifiers and different features on the original dataset. Top: 1k images, Bottom: 10k images.
-Figure 7: Average AUC for our approach considering different classifiers and different features on the dataset after random compression and resizing. Top: 1k images, Bottom: 10k images.
-Appendix A Additional ablation study
-In Section 4 of the paper we describe the CLIP-based detection procedure used in all experiments.
-Two important design choices outlined in the paper are that i) we extract features from the next to last layer of the CLIP ViT L/14 architecture; and ii) we use a SVM classifier, designed based on a limited set of reference features.
-Here we test some alternative solutions, that is:
+To this end a ResNet-50 is trained in two different ways, using ADM images (DIRE 1) and StyleGAN images (DIRE 2), respectively.
+  10. (j)
+    NPR [71] works on residual images computed as the difference between the original image and its interpolated version. The classifier is a ResNet-50 trained on only 4 classes of the ProGAN dataset.
+Refer to captionRefer to caption
+Refer to captionRefer to caption
+Figure 7: Average AUC for our approach considering different classifiers and different features on the original dataset (top) and after random compression and resizing (bottom).
+Left: 1k images, Right: 10k images.
+Appendix B Additional ablation study
+In Section 4 of the paper we describe the CLIP-based detector used in all experiments.
+In the proposed procedure two key design choices are made: i) we extract features from the next to last layer of the CLIP ViT L/14 architecture; and ii) we use a SVM classifier, trained on a limited set of reference features.
+Here we test alternative solutions, that is:
   • extracting features from the last layer of the architecture;
   • using other classifiers: Logistic Regression (LR), Mahalanobis distance (MAH), Gaussian Naive Bayes classifier (GNB), Soft voting k-Nearest Neighbor (SNN) [55].
-In Figures 6 and 7, we show the results in terms of average AUC over the dataset described in Section 3 (main paper) using both original images and images impaired by common post-processed steps such as recompression and resizing.
-For all detectors, we consider the versions 1k and 10k (1000 and 10000 reference images, respectively).
+In Figure 7, we show the results in terms of average AUC over the dataset described in Section 3 (main paper) using both original images and images impaired by common post-processing steps such as recompression and resizing.
+For all detectors, we consider the versions 1k and 10k (1000 and 10000 reference images per class, respectively).
 In all cases, the SVM classifier seems to ensure the best performance, followed closely by the Logistic Regression, while the other classifiers provide a less consistent performance.
@@
 These results motivated our design choices.
+Families of Generators
+GAN Diffusion Comm. Tools Average
+Method AUC/Acc AUC/Acc AUC/Acc AUC/Acc
+Wang et al. 92.1 / 69.4 55.8 / 50.3 50.2 / 49.8 66.0 / 56.5
+PatchFor. 84.9 / 53.6 76.4 / 50.4 50.5 / 50.0 70.6 / 51.3
+Grag. et al. 95.8 / 90.1 70.8 / 57.0 41.8 / 43.6 69.5 / 63.6
+Mand. et al. 88.9 / 81.6 55.8 / 53.7 22.6 / 32.9 55.7 / 56.1
+Liu et al. 99.0 / 89.7 79.9 / 69.5 30.8 / 42.9 69.9 / 67.4
+Corvi et al. 72.7 / 52.1 91.5 / 75.1 82.1 / 62.8 82.1 / 63.3
+LGrad 87.8 / 76.4 69.2 / 60.4 48.3 / 52.1 68.4 / 63.0
+Ojha et al. 96.1 / 85.4 82.0 / 63.4 73.8 / 68.3 84.0 / 72.4
+DIRE-1 65.9 / 64.4 71.6 / 72.8 50.2 / 49.9 62.5 / 62.4
+DIRE-2 65.1 / 60.1 70.4 / 65.8 45.3 / 49.9 60.3 / 58.6
+NPR 89.5 / 79.7 82.0 / 68.1 49.3 / 50.1 73.6 / 66.0
+Ours 1k 91.2 / 76.4 92.1 / 76.2 76.6 / 54.5 86.6 / 69.0
+Ours 1k+ 86.2 / 74.2 89.9 / 79.9 85.3 / 72.6 87.1 / 75.6
+Ours 10k 92.4 / 79.1 92.6 / 73.3 80.5 / 52.6 88.5 / 68.3
+Ours 10k+ 89.3 / 74.9 91.8 / 77.2 87.0 / 67.3 89.4 / 73.1
+Table 8: Comparison with SOTA methods in terms of AUC and Accuracy. We report the mean AUC and Accuracy for each family of generators and the global average.
+Families of Generators
+GAN Diffusion Comm. Tools Average
+Method AUC/Acc AUC/Acc AUC/Acc AUC/Acc
+Wang et al. 79.2 / 62.0 59.3 / 50.4 44.1 / 49.9 60.9 / 54.1
+PatchFor. 54.2 / 50.0 64.8 / 50.3 62.9 / 50.1 60.7 / 50.1
+Grag. et al. 89.0 / 67.6 67.9 / 50.8 47.3 / 50.0 68.1 / 56.1
+Mand. et al. 69.2 / 55.1 50.6 / 50.8 44.7 / 49.6 54.9 / 51.9
+Liu et al. 54.4 / 51.5 56.3 / 51.1 53.0 / 50.7 54.6 / 51.1
+Corvi et al. 74.0 / 55.1 77.3 / 62.1 52.1 / 50.1 67.8 / 55.8
+LGrad 52.0 / 50.9 48.0 / 50.5 49.4 / 50.4 49.8 / 50.6
+Ojha et al. 86.1 / 73.7 73.7 / 59.8 49.6 / 50.6 69.8 / 61.4
+DIRE-1 47.6 / 50.2 51.0 / 50.5 50.1 / 49.9 49.6 / 50.2
+DIRE-2 47.7 / 49.6 52.3 / 52.5 45.3 / 49.9 48.4 / 50.7
+NPR 48.7 / 50.2 50.0 / 49.8 56.0 / 50.1 51.6 / 50.0
+Ours 1k 76.2 / 68.9 80.6 / 71.2 72.2 / 60.7 76.3 / 66.9
+Ours 1k+ 75.0 / 67.2 87.8 / 77.8 83.4 / 65.7 82.1 / 70.3
+Ours 10k 76.9 / 63.6 81.1 / 63.7 73.0 / 52.4 77.0 / 59.9
+Ours 10k+ 78.2 / 69.0 89.3 / 78.7 84.7 / 66.4 84.1 / 71.4
+Table 9: AUC/Accuracy in the presence of post-processing. All images have been randomly cropped, resized and compressed to simulate a realistic scenario on the web.
+Appendix C AUC vs Accuracy
+In this Section we present additional results on the synthetic generators analyzed in the main paper aggregated by family (GAN, Diffusion, Commercial tools).
+Results are not only in terms of AUC but also of balanced accuracy computed with a fixed threshold of 0.5.
+Indeed in a practical scenario a threshold must be set to discriminate between real and fake images, which is not a trivial task as can be seen in Table 8.
+In fact, in the absence of an adequate calibration procedure, a significant drop between AUC and accuracy is often observed.
+However, our approach provides good results even in this challenging context, especially on post-processed images (Tab.
+9) where most methods are equivalent to flipping coins.
+Refer to captionRefer to captionRefer to caption
 Figure 8: Robustness analysis in terms of AUC carried out on the Stable Diff. 2, SDXL, DALL·E 3, DALL·E 2, Midjourney, and Firefly generation models of the SynthBuster dataset [4].
-Appendix B Additional robustness analysis
-In Section 5 of the paper, in Tab.
-6, we provide some results on the robustness of SoTA and proposed methods in the presence of image impairments.
-Due to lack of space, however, we show only some very synthetic results, averaged on generic post-processed images.
-Here, we analyze robustness in more detail as a function of the JPEG Quality Factor, which ranges from 100 to 60, and the resizing scale, which ranges from 125% to 25%.
-In addition, we consider also another compression method, WEBP, which is gaining popularity on social networks and has been rarely been used for augmentation.
+Appendix D Additional robustness analysis
+In Section 5 of the main paper we provide some results on the robustness of SoTA and proposed methods in the presence of image impairments.
+Due to lack of space, however, we show only some very synthetic results, averaged on all kinds of post-processed images.
+In Fig.
+8, we analyze robustness in more detail as a function of JPEG Quality Factor, ranging from 100 to 60, and resizing scale, going from 125% to 25%.
+In addition, we consider also another compression method, WebP, which is gaining popularity on social networks but has been rarely considered in experimental analyses.
 Images are from the SynthBuster dataset [4], generated by Stable Diffusion 2, SDXL, DALL·E 2, DALL·E 3, Midjourney, and Firefly.
-We show results (AUC) only for the most competitive (in terms of robustness) SoTA methods, Corvi and Ojha.
-For our method we consider again four variants similarly to the main paper: the versions using 1k or 10k real and fake reference images, and the corresponding versions, 1k+ and 10k+, where data have been augmented with randomly compressed/resized images, while keeping constant the total number of images.
+We show results (AUC) only for the SoTA methods that proved most competitive in terms of robustness, Corvi and Ojha.
+For our method we consider again the versions with 1k or 10k real and fake reference images, and the corresponding versions, 1k+ and 10k+, with the same number of images including random compression and resizing.
 Ojha suffers some performance losses, up to 10 points, in the presence of compression, be it JPEG or WebP, and a much more rapid decline with resizing, to the point of becoming useless in the presence of a 25% rescaling.
@@
 The most remarkable outcome of this experiment, however, is the impressive robustness of the CLIP-based detectors.
-Both the 1k+ and 10k+ versions, the ones with augmentation, are basically insensitive to compression, no matter JPEG or WEBP, and resizing.
+Both the 1k+ and 10k+ versions, those with augmentation, are basically insensitive to compression, no matter JPEG or WebP, and resizing.
 The versions without augmentation suffer some loss of performance but not nearly as dramatic as for the reference methods, and remain effective in all conditions.
+AUC/Acc Social network: X
+Method DALL·E3 Midjourney FireFly Average
+Wang et al. 22.5 / 50.0 36.7 / 50.0 65.3 / 50.0 41.5 / 50.0
+PatchFor. 54.5 / 50.0 59.9 / 50.0 46.5 / 50.0 53.6 / 50.0
+Grag. et al. 44.7 / 49.6 58.4 / 49.7 80.5 / 49.9 61.2 / 49.7
+Mand. et al. 42.9 / 49.3 53.7 / 50.6 52.7 / 50.1 49.8 / 50.0
+Liu et al. 34.5 / 49.7 38.7 / 49.7 54.9 / 49.9 42.7 / 49.8
+Corvi et al. 89.3 / 86.8 98.3 / 90.8 79.1 / 52.1 88.9 / 76.6
+LGrad 33.7 / 39.0 37.9 / 41.4 61.4 / 58.4 44.3 / 46.3
+Ojha et al. 34.4 / 45.6 32.2 / 46.2 73.0 / 65.7 46.5 / 52.5
+DIRE-1 54.3 / 52.0 60.7 / 54.3 55.1 / 51.1 56.7 / 52.5
+DIRE-2 60.0 / 55.0 58.6 / 53.9 47.7 / 46.5 55.4 / 51.8
+NPR 43.8 / 41.7 46.2 / 44.9 65.4 / 62.0 51.8 / 49.5
+Ours 1k 69.4 / 50.9 70.2 / 53.3 69.3 / 51.4 69.6 / 51.9
+Ours 1k+ 82.1 / 73.3 82.8 / 74.5 83.8 / 74.9 82.9 / 74.3
+Ours 10k 65.8 / 50.0 69.3 / 51.7 70.7 / 50.9 68.6 / 50.9
+Ours 10k+ 78.7 / 63.7 82.2 / 69.6 82.7 / 69.6 81.2 / 67.6
+Ours fusion 90.0 / 85.9 97.5 / 93.8 77.8 / 57.0 88.4 / 78.9
+Ours fusion+ 94.7 / 87.3 97.9 / 92.1 85.4 / 74.2 92.6 / 84.5
+Ours Few-Shot 98.5 / 92.5 95.8 / 87.4 92.5 / 83.5 95.6 / 87.8
+Table 10: Results in terms of AUC and accuracy of proposal and SOTA methods on real and synthetic images download from X.
+The last row reports a few-shot experiment, where we suppose to know in advance 10+10 real/fake images from a specific generator.
+Refer to caption Refer to caption Refer to caption Refer to caption
+Figure 9: Some examples of images downloaded from the social network X. From left to right: a real image, synthetic images from DALL·E 3 [6], Firefly [31], Midjourney [53].
+Appendix E Few-shot analysis in the wild
+In this section we present an experiment in a few-shot scenario.
+The idea is to explore the ability of our detector to work with very limited data and adapt to a situation where only a few real and synthetic examples are available.
+We downloaded 500 real images and 1,500 synthetic images from X from three different generators DALL·E 3, Midjourney and Firefly (some examples can be seen in Fig.
+9).
+To understand which generator was used to create a specific image, we relied on tags and annotations present on the social network.
+In our few-shot scenario, we take $10$ real images and $10$ generated images as examples from a specific model and test on all the others images (note that we present average results on 1,000 runs).
+Results are reported in Tab.
+10 in terms of AUC and accuracy.
+The availability of just 10+10 images of the target data provides an impressive performance boost with respect to the same method trained on a dataset much larger but not aligned with the test data (real images come from COCO, synthetic images from Latent diffusion models).
+It is important to underline that this is a realistic scenario, in which one is called to decide on images generated by a new method and a few sample images are available as a support.
+We believe that this can represent an interesting direction for the application in the wild or to easily adapt a detector to more challenging situations where some prior information is available.
+To complete our analysis on this dataset we include in Tab.
+10 the comparison with SoTA methods and our original proposal.
+We can observe that the performance of all the methods degrade which highlights the increased difficulty to handle a realistic scenario over the web.
+The best performance can be obtained by using our fusion approach that take the best of the low-level and high-level features and is able to achieve on average more than 90% in terms of AUC which is comparable with the few-shot analysis where some prior knowledge is available.
 Generator
@@
 The last column reports the resolution of the images in the dataset.
-Appendix C Further generalization results
-In this Section we extend our generalization analysis to additional data.
-In the main paper we included 6,000 real images and 18,$0$ fake images from 18 models, now we consider 8 more generators with 8,000 additional synthetic images for our test set (Tab.
-7).
-The StyleGAN-T [63], eDiff-I [3], and GigaGAN [39] images are provided by the authors of the original papers.
-Instead, we ourselves generated images for Score-SDE [68], Stable Diff.
-[62], DiT [57], Deepfloyd-IF [44], Diff.ProjectedGAN [73], GALIP [70], and DDPM [34] using the pre-trained models available on-line.
-The other generated images come from three public datasets [71, 14, 4].
-To avoid possible biases, performance metrics are evaluated using for each fake dataset images drawn from the real dataset used in training, indicated in the central columns of Table 1 of main paper and Table 7 of this supplemental.
-AUC/Acc BigGAN EG3D
-Diff. Proj.
-GALIP
-Taming
-DALL·E
-DDPM
-Deepfloyd-IF
-AVG
-Method
-GAN
-Transf.
-Mini
-II stage
-[71] Wang et al.
-92.7 / 66.1
-94.4 / 59.2
-89.8 / 52.1
-89.7 / 57.4
-54.3 / 51.7
-62.5 / 51.8
-31.6 / 50.1
-43.1 / 50.1
-69.8 / 54.8
-[9] PatchFor.
-85.5 / 52.5
-69.8 / 50.0
-92.6 / 61.7
-98.2 / 73.4
-71.2 / 51.0
-83.8 / 51.4
-98.4 / 50.2
-83.4 / 50.0
-85.4 / 55.0
-[33] Grag. et al.
-98.7 / 94.2
-98.9 / 93.8
-100. / 72.5
-96.2 / 79.5
-90.3 / 76.4
-83.4 / 62.5
-49.7 / 45.1
-71.2 / 61.3
-86.0 / 73.2
-[47] Liu et al.
-94.7 / 81.3
-99.0 / 86.3
-99.5 / 84.9
-94.3 / 78.4
-95.4 / 78.9
-98.4 / 88.1
-22.8 / 49.6
-97.4 / 86.8
-87.7 / 79.3
-[14] Corvi et al.
-83.4 / 51.8
-25.2 / 50.0
-96.6 / 71.2
-87.7 / 50.9
-99.3 / 89.5
-99.7 / 83.8
-100. / 90.3
-68.5 / 50.8
-82.5 / 67.3
-[69] LGrad
-77.2 / 69.0
-68.8 / 59.8
-99.5 / 90.1
-56.7 / 55.1
-74.1 / 64.5
-67.3 / 59.9
-9.8 / 16.4
-75.0 / 62.5
-66.1 / 59.7
-[56] Ojha et al.
-99.6 / 96.4
-92.6 / 82.5
-97.4 / 75.2
-98.6 / 89.9
-94.1 / 84.8
-97.1 / 84.9
-77.7 / 68.2
-60.8 / 50.0
-89.7 / 79.0
-[72] DIRE-1
-99.8 / 95.3
-50.1 / 50.0
-49.8 / 51.6
-100. / 96.7
-73.1 / 72.4
-99.7 / 96.5
-23.1 / 50.0
-99.4 / 95.8
-74.4 / 76.0
-[72] DIRE-2
-98.6 / 82.4
-46.1 / 50.0
-50.2 / 51.4
-99.3 / 81.8
-77.0 / 66.2
-98.7 / 81.8
-14.0 / 49.8
-95.6 / 80.5
-72.4 / 68.0
-Ours 1k
-96.7 / 85.2
-86.3 / 73.8
-99.3 / 69.7
-100. / 99.7
-99.3 / 94.3
-100. / 98.9
-94.7 / 86.0
-99.7 / 90.6
-97.0 / 87.3
-Ours 1k+
-90.9 / 77.4
-75.5 / 64.7
-89.4 / 68.4
-99.9 / 98.2
-93.4 / 82.3
-99.6 / 94.0
-94.6 / 87.7
-99.7 / 95.1
-92.9 / 83.5
-Ours 10k
-98.2 / 85.9
-90.8 / 81.4
-99.8 / 90.1
-100. / 99.9
-99.6 / 94.3
-100. / 99.1
-94.9 / 88.3
-99.8 / 88.1
-97.9 / 90.9
-Ours 10k+
-91.7 / 75.0
-80.7 / 72.5
-88.9 / 74.6
-100. / 97.8
-92.5 / 77.2
-99.8 / 93.5
-95.9 / 79.5
-99.8 / 92.8
-93.7 / 82.8
+AUC/Acc BigGAN EG3D Diff. Proj. GALIP Taming DALL·E DDPM Deepfloyd-IF AVG
+Method GAN Transf. Mini II stage
+Wang et al. 92.7 / 66.1 94.4 / 59.2 89.8 / 52.1 89.7 / 57.4 54.3 / 51.7 62.5 / 51.8 31.6 / 50.1 43.1 / 50.1 69.8 / 54.8
+PatchFor. 85.5 / 52.5 69.8 / 50.0 92.6 / 61.7 98.2 / 73.4 71.2 / 51.0 83.8 / 51.4 98.4 / 50.2 83.4 / 50.0 85.4 / 55.0
+Grag. et al. 98.7 / 94.2 98.9 / 93.8 100. / 72.5 96.2 / 79.5 90.3 / 76.4 83.4 / 62.5 49.7 / 45.1 71.2 / 61.3 86.0 / 73.2
+Mand. et al. 92.2 / 83.0 100. / 99.8 64.5 / 49.9 77.6 / 59.1 91.8 / 84.1 83.6 / 69.8 99.9 / 97.4 49.2 / 48.9 82.3 / 74.0
+Liu et al. 94.7 / 81.3 99.0 / 86.3 99.5 / 84.9 94.3 / 78.4 95.4 / 78.9 98.4 / 88.1 22.8 / 49.6 97.4 / 86.8 87.7 / 79.3
+Corvi et al. 83.4 / 51.8 25.2 / 50.0 96.6 / 71.2 87.7 / 50.9 99.3 / 89.5 99.7 / 83.8 100. / 90.3 68.5 / 50.8 82.5 / 67.3
+LGrad 77.2 / 69.0 68.8 / 59.8 99.5 / 90.1 56.7 / 55.1 74.1 / 64.5 67.3 / 59.9  9.8 / 16.4 75.0 / 62.5 66.1 / 59.7
+Ojha et al. 99.6 / 96.4 92.6 / 82.5 97.4 / 75.2 98.6 / 89.9 94.1 / 84.8 97.1 / 84.9 77.7 / 68.2 60.8 / 50.0 89.7 / 79.0
+DIRE-1 99.8 / 95.3 50.1 / 50.0 49.8 / 51.6 100. / 96.7 73.1 / 72.4 99.7 / 96.5 23.1 / 50.0 99.4 / 95.8 74.4 / 76.0
+DIRE-2 98.6 / 82.4 46.1 / 50.0 50.2 / 51.4 99.3 / 81.8 77.0 / 66.2 98.7 / 81.8 14.0 / 49.8 95.6 / 80.5 72.4 / 68.0
+NPR 86.8 / 77.2 53.3 / 57.5 100. / 99.2 90.7 / 77.6 80.2 / 69.2 79.0 / 73.3 92.4 / 61.6 90.9 / 76.8 84.2 / 74.1
+Ours 1k 96.9 / 86.1 87.0 / 73.0 99.3 / 70.2 100. / 99.7 99.4 / 95.0 100. / 99.1 95.1 / 86.7 99.7 / 90.7 97.2 / 87.6
+Ours 1k+ 92.0 / 80.0 76.4 / 65.5 89.6 / 67.7 99.9 / 98.6 93.5 / 83.0 99.6 / 95.0 94.9 / 87.4 99.7 / 96.0 93.2 / 84.1
+Ours 10k 98.2 / 87.0 87.7 / 63.2 99.7 / 87.4 100. / 99.8 99.7 / 97.4 100. / 99.4 93.9 / 61.9 99.8 / 85.9 97.4 / 85.3
+Ours 10k+ 93.1 / 77.5 79.9 / 58.8 92.5 / 76.0 100. / 97.0 94.9 / 79.5 99.9 / 94.4 98.1 / 91.5 99.8 / 92.2 94.8 / 83.3
 Table 12: Comparison with SOTA methods on additional data.
@@
 For our approach we show four variants: trained on 1k real and 1k fake images; 10k real and 10k fake images; trained on 1k and 1k fake images but including compressed/resized images (1k+) and trained on 10k and 10k fake images but including compressed/resized images (10k+).
-AUC/Acc BigGAN EG3D
-Diff. Proj.
-GALIP
-Taming
-DALL·E
-DDPM
-Deepfloyd-IF
-AVG
-Method
-GAN
-Transf.
-Mini
-II stage
-[71] Wang et al.
-82.5 / 55.5
-84.7 / 52.2
-80.8 / 53.2
-92.2 / 59.2
-66.2 / 50.6
-66.7 / 50.5
-69.6 / 50.2
-47.9 / 49.9
-73.8 / 52.7
-[9] PatchFor.
-58.7 / 50.3
-60.3 / 50.0
-55.6 / 50.0
-72.3 / 50.3
-50.4 / 50.1
-71.0 / 50.1
-82.9 / 51.0
-69.7 / 50.0
-65.1 / 50.2
-[33] Grag. et al.
-97.6 / 74.4
-92.7 / 56.5
-99.2 / 65.6
-99.1 / 79.8
-83.6 / 53.3
-89.6 / 54.0
-74.5 / 49.4
-54.3 / 50.0
-86.3 / 60.4
-[47] Liu et al.
-57.7 / 52.0
-51.6 / 50.4
-62.9 / 54.1
-65.2 / 52.3
-49.2 / 50.2
-49.1 / 50.7
-58.3 / 49.1
-64.4 / 52.0
-57.3 / 51.4
-[14] Corvi et al.
-77.3 / 53.0
-64.5 / 59.7
-93.3 / 59.8
-87.4 / 51.5
-94.4 / 76.7
-97.4 / 74.8
-74.7 / 59.6
-70.4 / 50.9
-82.4 / 60.7
-[69] LGrad
-53.5 / 51.6
-53.2 / 51.1
-56.9 / 51.1
-57.0 / 50.3
-51.3 / 51.1
-41.7 / 49.1
-56.0 / 49.7
-38.7 / 48.8
-51.0 / 50.4
-[56] Ojha et al.
-94.8 / 84.9
-79.3 / 61.5
-84.7 / 68.5
-95.7 / 84.5
-92.9 / 81.8
-89.1 / 72.8
-90.8 / 79.1
-59.1 / 49.5
-85.8 / 72.8
-[72] DIRE-1
-47.2 / 49.5
-35.2 / 49.3
-35.2 / 49.6
-47.4 / 50.6
-47.3 / 49.7
-54.6 / 51.1
-38.5 / 49.4
-64.6 / 53.2
-46.2 / 50.3
-[72] DIRE-2
-44.6 / 47.0
-33.6 / 47.1
-37.5 / 44.9
-46.2 / 47.2
-43.3 / 46.8
-53.1 / 50.0
-42.2 / 47.5
-66.3 / 59.3
-45.9 / 48.7
-Ours 1k
-84.0 / 74.8
-64.4 / 59.5
-75.3 / 68.3
-98.5 / 93.0
-85.3 / 75.0
-96.5 / 89.0
-76.7 / 66.6
-90.6 / 80.1
-83.9 / 75.8
-Ours 1k+
-88.9 / 78.2
-60.0 / 54.6
-70.5 / 60.9
-99.8 / 98.3
-86.5 / 77.3
-98.9 / 93.6
-96.1 / 84.7
-99.4 / 96.2
-87.5 / 80.5
-Ours 10k
-84.4 / 59.1
-70.9 / 63.3
-68.8 / 55.6
-99.2 / 79.7
-85.7 / 67.7
-97.5 / 68.7
-64.6 / 57.0
-95.1 / 62.2
-83.3 / 64.1
-Ours 10k+
-91.3 / 79.4
-68.7 / 59.7
-68.7 / 61.7
-99.9 / 99.0
-86.2 / 71.2
-99.4 / 93.6
-92.7 / 62.4
-99.8 / 97.8
-88.3 / 78.1
+AUC/Acc BigGAN EG3D Diff. Proj. GALIP Taming DALL·E DDPM Deepfloyd-IF AVG
+Method GAN Transf. Mini II stage
+Wang et al. 82.5 / 55.5 84.7 / 52.2 80.8 / 53.2 92.2 / 59.2 66.2 / 50.6 66.7 / 50.5 69.6 / 50.2 47.9 / 49.9 73.8 / 52.7
+PatchFor. 58.7 / 50.3 60.3 / 50.0 55.6 / 50.0 72.3 / 50.3 50.4 / 50.1 71.0 / 50.1 82.9 / 51.0 69.7 / 50.0 65.1 / 50.2
+Grag. et al. 97.6 / 74.4 92.7 / 56.5 99.2 / 65.6 99.1 / 79.8 83.6 / 53.3 89.6 / 54.0 74.5 / 49.4 54.3 / 50.0 86.3 / 60.4
+Mand. et al. 71.4 / 55.5 85.8 / 62.1 60.3 / 49.6 78.1 / 57.3 82.3 / 59.7 68.2 / 57.4 61.1 / 58.6 49.0 / 49.8 69.5 / 56.3
+Liu et al. 57.7 / 52.0 51.6 / 50.4 62.9 / 54.1 65.2 / 52.3 49.2 / 50.2 49.1 / 50.7 58.3 / 49.1 64.4 / 52.0 57.3 / 51.4
+Corvi et al. 77.3 / 53.0 64.5 / 59.7 93.3 / 59.8 87.4 / 51.5 94.4 / 76.7 97.4 / 74.8 74.7 / 59.6 70.4 / 50.9 82.4 / 60.7
+LGrad 53.5 / 51.6 53.2 / 51.1 56.9 / 51.1 57.0 / 50.3 51.3 / 51.1 41.7 / 49.1 56.0 / 49.7 38.7 / 48.8 51.0 / 50.4
+Ojha et al. 94.8 / 84.9 79.3 / 61.5 84.7 / 68.5 95.7 / 84.5 92.9 / 81.8 89.1 / 72.8 90.8 / 79.1 59.1 / 49.5 85.8 / 72.8
+DIRE-1 47.2 / 49.5 35.2 / 49.3 35.2 / 49.6 47.4 / 50.6 47.3 / 49.7 54.6 / 51.1 38.5 / 49.4 64.6 / 53.2 46.2 / 50.3
+DIRE-2 44.6 / 47.0 33.6 / 47.1 37.5 / 44.9 46.2 / 47.2 43.3 / 46.8 53.1 / 50.0 42.2 / 47.5 66.3 / 59.3 45.9 / 48.7
+NPR 52.6 / 50.0 40.2 / 49.8 48.8 / 50.3 57.8 / 50.6 46.2 / 49.8 51.7 / 49.6 54.5 / 50.5 51.3 / 48.9 50.4 / 49.9
+Ours 1k 84.6 / 75.8 64.8 / 60.3 73.4 / 66.7 98.3 / 92.6 85.1 / 75.8 96.2 / 88.9 76.1 / 67.7 90.0 / 79.9 83.6 / 76.0
+Ours 1k+ 89.6 / 79.7 60.5 / 55.1 69.6 / 61.0 99.9 / 98.4 86.9 / 78.0 98.9 / 93.7 96.4 / 84.8 99.5 / 96.5 87.7 / 80.9
+Ours 10k 84.1 / 67.0 64.3 / 54.9 73.7 / 62.9 98.1 / 83.9 86.2 / 74.0 96.8 / 77.8 70.9 / 60.7 91.3 / 67.5 83.1 / 68.6
+Ours 10k+ 92.4 / 82.4 63.1 / 55.5 74.1 / 63.9 99.9 / 98.9 88.1 / 78.2 99.4 / 94.9 96.1 / 87.1 99.7 / 97.0 89.1 / 82.2
 Table 13: Comparison with SOTA methods on additional data in the presence of post-processing.
@@
 The entries in bold underline the best performance for each dataset.
-In Tab.
-8 we show the results in terms of AUC and Accuracy on such data for the four version of the proposed CLIP-based detector and for the SoTA methods described in the main paper (Section 5): Wang et al.
-[71], Grag et al.
-[33], PatchForensics [9], Liu et al.
-[47], LGrad [69], Corvi et al.
-[14], Ojha et al.
-[56], DIRE [72].
+Appendix F Further generalization results
 […]
-Appendix D Attack scenario
-In this section, we conduct some additional experiments on a specific attack scenario, where a malicious user takes real images and passes them through the autoencoder of a specific generative model.
-The objective is to insert the low-level traces typical of synthetic architectures in real images in order to fool forensic detectors and make everything appear as synthetic.
-To this end, we took 1000 real images from the RAISE dataset [16] and passed them through the autoencoders of three different versions of diffusion models: Stable Diffusion 1.4, Stable Diffusion 2.0 [62] and Stable Diffusion XL [58].
-In Fig.
-9 we show that such process insert exactly the same forensic clues in the real images and they appear indistinguishable from the generated ones, without altering their visual appearance (see Fig.
-5 of the main paper).
-Refer to caption Refer to caption Refer to caption
-Refer to caption Refer to caption Refer to caption
-Figure 9: Top (from left to right): Fourier spectra of the noise residuals for synthetic images generated with Stable Diffusion 1.4, Stable Diffusion 2.0 and Stable Diffusion XL.
-Bottom: Fourier spectra of the noise residuals for the real images that passed through the corresponding autoencoder of the synthetic generators.
-We can observe that this operation has inserted in the real images exactly the same artifacts present in the synthetic ones.
-In Table 10 we report results in terms of accuracy before and after the attack for our method and the best competitors.
-Before the attack, the detector must discriminate between real and generated images while, after the attack, it must discriminate between laundered real images and generated ones.
-We notice that before the attack the method proposed by Corvi et al.
-can achieve almost perfect results.
-However, it is very sensitive to the insertion of low-level traces and provides totally wrong results on the attacked data.
-For Ojha et al the performance also worsen, with an average 7% loss.
-Instead, for our method the Accuracy is almost unaffected by the attack, with an average loss below 1.3%.
-Based on these results, it appears that the proposed CLIP-based detector does not rely on the same low-level traces exploited by most of the current detectors.
-This ensures higher robustness to common attacks and opens the way to the use of suitable fusion strategies.
-However, there is much room for further analysis, to begin with identifying which traces our detector is actually exploiting.
-Accuracy before the attack after the attack
-Method
-St.Diff.
-St.Diff.
-SDXL
-St.Diff.
-St.Diff.
-SDXL
-1.4
-2.0
-1.4
-2.0
-Corvi et al.
-100.
-99.2
-99.7
-50.0
-49.2
-53.7
-Ojha et al.
-69.9
-75.6
-68.0
-59.0
-69.0
-64.5
-Ours 1k
-63.6
-64.3
-60.2
-63.6
-63.8
-59.2
-Ours 1k+
-76.8
-77.0
-75.2
-77.0
-75.5
-72.6
-Ours 10k
-65.8
-60.5
-54.4
-65.9
-60.2
-54.0
-Ours 10k+
-78.8
-72.8
-70.0
-79.1
-71.5
-68.7
-Table 10: Results in terms of Accuracy before and after the laundering attack on real images.
-Before the attack all methods achieve a very good performance.
-After the attack the performance of our method is almost unchanged, while significant losses are observed for the best reference methods, especially for Corvi et al.
-that goes from perfect detection to random guess.
-Appendix E Few-shot analysis in the wild
-In this section we present an experiment in a few-shot scenario.
-The idea is to explore the ability of our detector to work with very limited data and adapt to a situation where only a few real and synthetic examples are available.
-We downloaded real and synthetic images from Twitter and considered three different generators DALL·E 3, Midjourney and Firefly (some examples can be seen in Fig.
-10).
-To understand which generator was used to create a specific image, we relied on tags and annotations present on the social network.
-In our few-shot scenario, we take $10$ real images and $10$ generated images as examples from a specific model and test on all the others images (note that we present average results on 1000 runs).
-Results are reported in Tab.
-11 in terms of AUC and Accuracy.
-The availability of just 10+10 images of the target data provides an impressive performance boost with respect to the same method trained on a larger dataset, but not aligned with the test data (real images come from COCO, synthetic images from Latent diffusion models).
-It is important to underline that this is a realistic scenario, in which one is called to decide on images generated by a new method and a few sample images are available as a support.
-We believe that this can represent an interesting direction for the application in the wild or to easily adapt a detector to more challenging situations where some prior information is available.
-Refer to caption Refer to caption Refer to caption Refer to caption
-Figure 10: Some examples of images downloaded from the social network Twitter. From left to right: a real image, synthetic images from DALL·E 3 [6], Firefly [30], Midjourney [52].
-AUC/Acc Twitter
-Method
-DALL·E3
-Midjourney
-FireFly
-Average
-Ours 1k
-71.9 / 51.1
-70.3 / 54.0
-77.4 / 52.4
-73.2 / 52.5
-Ours 1k+
-83.6 / 72.3
-75.5 / 67.3
-89.6 / 80.5
-82.9 / 73.4
-Ours 10k
-68.2 / 50.5
-72.0 / 51.5
-70.7 / 51.5
-70.3 / 51.2
-Ours 10k+
-67.0 / 60.0
-78.3 / 71.0
-74.3 / 68.5
-73.2 / 66.5
-Ours Few-Shot
-97.7 / 90.6
-91.9 / 82.5
-96.6 / 88.4
-95.4 / 87.2
-Table 11: We show results in a few-shot experiment, where we suppose to know in advance 10+10 real/fake images from a specific generator and compare them with our basic approach trained on a larger set of real/fake images from misaligned data.
-As can be seen, even with just 10 images it is possible to adapt the model and obtain a large improvement.